### v1

# Benchmarking General Purpose In-Context Learning

Abstract
In-context learning (ICL) capabilities is becoming increasingly appealing towards building general intelligence. Taking this concept one step further, we draw a parallel to humans and many animals, who inherit primarily learning capabilities but refine their memory and acquire diverse skills and knowledge through extensive lifelong experiences. This parallel inspires our approach to general purpose in-context learning (GPICL). This paper introduces two lightweight but insightful benchmarks specifically crafted to train and evaluate GPICL functionalities. Each benchmark encompasses a wide range of diverse tasks characterized by generation and interaction, minimal transferable knowledge, and long-term dependency. These features present significant challenges for models that primarily rely on context or interactions to enhance their proficiency. We hope that these benchmarks will not only advance research in GPICL but also contribute significantly to the broader field of general intelligence.
00footnotetext: Works done at Baidu Inc. and University of Science and Technology of China.00footnotetext: Corresponds to: fanwang.px@gmail.com and forrest@ustc.edu.cn00footnotetext: source code available at https://github.com/FutureAGI/L3C
1 Introduction
The success of large-scale generative language models can be primarily attributed to two key factors: 1. Zero-shot generalization derived from the extensive accumulation and storage of knowledge within the model parameters during the pretraining and fine-tuning stages (Kojima et al., 2022); 2. In-context learning(ICL) to distill knowledge from the context during the inference stage (Brown et al., 2020). The ability of zero-shot generalization remains unchanged post-training, which parallel the concept of “innate abilities” (Koulakov et al., 2021) of biology. The ICL capability allows the agent to handle various unseen tasks without modifying the model parameters, only resulting in changes to the hidden states and memories. Currently, the ICL capabilities of large language models are confined to tasks that solely involve interpreting natural language tokens and following instructions (i.e., mimicking demonstrations). However, they generally lack the ability to continually interact with real-world environments and engage in reinforcement learning from contexts. Additionally, extremely complex tasks that require substantial amounts of context still pose significant challenges.
Methodology Zero-shot Capability ICL
Potential
ICL
Horizon
Multi-task Learning Medium Low Short
Meta-Learning Low Medium Short
ICL High Medium Medium
GPICL Low High Long
Figure 1: Conceptual diagram of in-context learning (ICL) and general purpose in-context learning (GPICL)
The community has long recognized a significant gap between artificial intelligence and biological intelligence, particularly regarding innate abilities and the functionality to continuously learn throughout their lifespans (Parisi et al., 2019). For example, the initial limited capabilities of mammal infants stand in stark contrast to their skills in lifelong learning and adaptation. This disparity is further highlighted by the phenomenon known as the genomics bottleneck. Understanding this could guide the design of machine intelligence systems (Zador, 2019; Kirsch & Schmidhuber, 2021; Wang et al., 2022). A promising research area is general purpose in-context learning (GPICL) (Kirsch et al., 2022; Wang et al., 2022). GPICL involves meta-training across a wide range of task classes and learning to interpret context within a variety of highly diverse tasks. However, the characteristic of diversity in this context can be challenging to quantify. We aim to enrich the concept of GPICL by incorporating additional features, as shown in Figure 2. Consider an agent that ideally utilizes all available data, including meta-training data and contextual information; The GPICL system should corresponds to low zero-shot capability, high ICL potential, and an extended ICL horizon.
(a) LLM mainstream learning pipeline
(b) GPICL learning pipeline
Figure 2: Learning pipelines based on GPICL versus the current LLM learning pipeline (Ouyang et al., 2022).
Why GPICL? The objective of extending the in-context learning (ICL) horizon and potential is widely acknowledged. However, it raises the question of why we should reduce zero-shot capability, which appears to contradict the prevailing trend in large language models. First, inspired by the genomics bottleneck and building on previous work (Kirsch & Schmidhuber, 2021; Wang et al., 2022), one potential strategy to enhance generalizability of the meta-model is to maintain compact meta-parameters that are focused primarily on the functionality of learning itself, rather than to keep overwhelming parameters designed to capture all task-specific knowledge and skills. Second, the use of abundant synthetic data in pre-training and meta-training, which tends to be non-factual and biased, presents challenges to the usage of the model. In this context, GPICL could offer an ideal solution: since the meta-training stage can result in the loss of much task-specific knowledge and skills but preserves the core learning functionality, it helps prevent the meta-model from being overly influenced by non-factual information and bias in those tasks. Third, in practice, a model’s performance typically depends on both in-context learning (ICL) and zero-shot generalization capabilities, making it challenging to investigate and improve ICL independently. Isolating the ICL capability alone would be beneficial for inspiring better model architectures and learning algorithms.
Building on those points, we may futher propose an prospective approach for advancing the development of general intelligence in contemporary settings: Initially, a substantial volume of synthesized data, potentially containing a wealth of nonfactual and biased information, is used exclusively to cultivate the GPICL functionality. Subsequently, a considerable amount of high-quality, unbiased, well-aligned data is employed to enrich the contexts and memory, enabling the agent to achieve capabilities approaching those of humans. A visual illustration of this proposed pipeline with existing methodologies is depicted in Figure 2.
Guidelines for designing GPICL benchmarks. We note that a major impediment to research in this area is the lack of benchmarks that adhere to the GPICL standard. We suggest that those benchmarks should satisfy the following standard:
  • Vast number of variant tasks with minimal knowledge transfer: To minimize the effectiveness of zero-shot generalization, it is advantageous to engage with a variety of tasks where task-specific knowledge cannot be transferred among different tasks. This guarantees that the model’s performance is primarily driven by its in-context learning capabilities rather than its ability of zero-shot generalization. It is essential to require numerous tasks to prevent the meta-model from hacking in-context learning (ICL) by merely memorizing all possible tasks and performing task identification, a limitation that most meta-learning benchmarks fall into when given limited tasks.
  • Interactive and generative tasks: GPICL should include task exploration and completion processes that are both interactive and generative, such that in-context reinforcement learning capabilities can be effectively meta-trained and evaluated. Consequently, benchmarks that solely focus on encoding a single sample characterized by extended contexts and features and conducting single-step predictions or classifications do not meet the criteria.
  • Lifelong in-context learning: GPICL should handle very long-term dependencies that can scale up to millions or even billions of steps, effectively supporting the learning process over extensive periods (ICL horizon $T$ should be “from cradle to grave”). It is necessary to clarify that $T$ should denote the “minimum steps” possible given “ideal ICL functionality”. While many previous gradient-based meta-learning methods also encode adaptation over millions of samples (e.g., many gradient-based meta-learning methods), most of these are to blame on the inefficient learning functionality. Consequently, they should not be considered as representing a long ICL horizon.
Contribution of this paper: We propose two critical benchmarks that adhere to the GPICL standard and are capable of generating unlimited synthetic data: 1. Meta-Language: We create a randomized yet consistent language pattern to meta-train models, demonstrating that a model, without prior exposure to natural language corpora, can acquire the ability to learn a new language from scratch through ICL; 2. Maze World: We generate an unlimited number of tasks within randomized mazes, where the model is meta-trained solely through context to learn exploration and navigation. For both benchmarks, we provide limited experiment results of baseline models based on Transformer (Vaswani et al., 2017) backbone. The primary goal of this paper is neither to introduce a new model or learning methodology, nor to provide an exhaustive survey of the existing ICL systems; rather, we aim to demonstrate through these baselines the potential for these benchmarks to benefit the research of ICL and GPICL.
2 Learning to learn an randomized language: Meta-Language
2.1 Problem Setting
Most large language models (LLMs) are trained on specific, existing languages. Rather than synthesizing existing languages, we primarily consider learning a new language as a task to focus more on the ICL functionality. Although there are hundreds of languages in the world, treating each language as a sample for GPICL provides only a few hundred samples, which is insufficient. We introduce the "meta language" benchmark. This benchmark is designed to generate a vast number of new "languages" and assess the models’ ability to learn an arbitrary new "language." This pseudo-language is simulated using n-gram settings, generated by a randomized neural network $f$. The probability of the next token is determined by its preceding n tokens and a set of random parameters $\theta$.
$p(x_{t}|x_{t-1},x_{t-2},...)=f(x_{t-1},...,x_{t-n};\theta),$ (1)
Algorithm 1 Generating a length=T sequence of meta langauge
1:Randomly Sample $\theta\sim\mathcal{N}(0,\sigma^{2})$, ($\theta$ is the parameter of the specified n-gram model $f$)
2:for t in [1, T] do
3:     $z_{1},...,z_{N}=f(x_{t-1},...,x_{t-n};\theta)$
4:     For $i\in[1,N]$, $z_{i}=\lambda[z_{i}-\mathbb{E}(z)]/\sqrt{\mathbb{V}(z)}$
5:     $p(x_{t}=x_{i})={exp(z_{i})}/{\sum_{j}exp(z_{j})}$
6:     Randomly sample $x_{t}\sim p(x_{t})$
The generation process is as simple as Algorithm 1. To maintain the diversity of the generated meta language, we also normalize the mean and variance of the logits input to Softmax to keep the distribution being neither too stiff nor too flat. We set the hyper-parameter $\lambda$ in Algorithm 1 such that the perplexity of the ground truth generator $\mathbb{E}[-logp(x_{t}))]$ is kept between $0.5$ and $1.0$. The sequences generated by the randomized generator are meaningless and chaotic. However, as long as the generator’s parameters remain static, a model can progressively learn and capture the underlying rules of the randomized generator. The complexity of the generated sequences can be bounded by $N^{n}$, where $N$ is the size of the vocabulary. Therefore, with a sufficiently large $N$ and $n$—parameters that can be adjusted as needed—the complexity of the pseudo-language can be made arbitrarily high to test super-long-term dependencies.
2.2 Baseline Solution
As the training data, we generated 500K sequences with a pre-defined vocabulary size of $N=32$, and the complexity of the pseudo-language was uniformly sampled across $n\in{3,4,5,6}$. Each sequence comprises $4,096$ tokens, totaling 2B tokens. We apply the auto-regressive transformer (Vaswani et al., 2017) with rotary embedding (Su et al., 2024). The hyper-parameter settings include: a tiny-sized transformer (303k parameters), a small-sized transformer (9.5m parameters), and a standard-sized(151m parameters) model. The evaluation set consists of independently sampled sequences, totaling 4K samples (16M tokens) for each $n\in\{2,3,4,5,6,7,8\}$, sampled uniformly. We assessed the model’s performance by plotting the averaged perplexity ($-\log p(x_{t})$) for each position across different $n$ values. To assess how the trained model generalizes, we also included the PG-19 (Rae et al., 2019) test set as a natural corpus. In this set, words are broken down into 26 alphabetic characters and randomly mapped onto 32 pre-defined tokens, the left slot are given to punctuation marks. The entire test texts are concatenated into a long sequence and broken down into sentences of 4,096 characters each. It is important to note that the model has never been trained on any natural language corpora. Specific details are provided in the appendices.
2.3 Evaluation Results
2.3.1 Towards Meta-Language Model
Figure 3: Perplexity per position on evaluation data sampled by different language complexity.
We demonstrate that this pseudo-language trained model can be regarded as a meta-language model designed to adapt to any conceivable language through GPICL. See from basic evaluation results of 151M model on the test sets in Figure 3, several observations are noteworthy: First, perplexity improves with increasing context length across all values of $n$, indicating that the model is effectively leveraging in-context learning. Notably, for $n=7$ and $n=8$—complexities on which the model was not explicitly trained—we observed a consistent semi-logarithmic improvement in perplexity with increasing context length. Second, a higher language complexity, indicated by larger values of $n$, correlates with a more gradual improvement in perplexity relative to context length. This phenomenon confirms that increased language complexity introduces longer-term dependencies within contexts, establishing meta language as an ideal benchmark for evaluating long-dependency modeling capabilities. Most importantly, the model demonstrates a significant performance gain through in-context learning (ICL) in realistic natural language scenarios, despite never being exposed to a single real-world natural word, illustrating that the ICL capability acquired through meta-language is not restricted to a specific language type.
We test the ability of generation of the meta-language model on two rudimentary real-world language tasks: English vocabulary learning and basic mathematical operations (addition and subtraction), which are typical for primary school education. For the English vocabulary task, the vocabulary consists of alphabet characters and common punctuation marks; for the math task, it includes digits and mathematical operators. The input contexts for testing included 40 English words and simple arithmetic operations below $5$. To discourage direct replication, we intentionally introduced errors into $15\%$ of the words or equations. We analyzed the outputs of the MetaLM across two different context lengths, with the shorter contexts comprising only the first half of the longer contexts. The results, depicted in Figure 4, show that MetaLM is capable of learning to memorize and correct both English vocabulary and mathematical equations, validating that its ICL capability is independent of vocabulary settings. Furthermore, the longer the context provided, the fewer errors appeared in the outputs, indicating a clear trend of improvement.
(a) English vocabulary learning task.
(b) Mathematical operation task.
Figure 4: Input contexts were designed such that $15\%$ contained errors. All mistakes are highlighted with red squares. The vocabulary for both tasks was limited to 32 pre-defined tokens, and there was no pre-training or further tuning applied to either task. Notice that this setup tests the MetaLM’s ability to adapt and correct based on in-context learning alone, without relying on prior knowledge or adjustments. The model under evaluation has not been trained on any natural langauge corpora. Additionally, the vocabulary is randomly assigned to the predefined 32 tokens prior to inference.
2.3.2 Scaling Laws for GPICL
We have conducted further research into how the scale of parameters affects GPICL capability. As illustrated in Figure 5, our findings reveal notable trends: 1. Language complexity ($n$) influences the relative performance of different-sized models and affects the dependence on context length. In simpler tasks ($n<4$), performance rapidly converges to lower bounds asymptotically ($t\sim T=4k$. Here, with a slight abuse of notation, $T$ does not necessarily represent the ICL horizon, but rather denotes the maximum context length used in our experiments.), and the performances of various models become indistinguishable, despite starting differently. 2. It should be noted that for tasks as straightforward as $n=2$, where a 300K model achieves comparable asymptotic performance to a 150M model, we observe that larger models cannot attain higher performance without sufficient context. It is reasonable to infer that increasing model complexity further will not be beneficial. This observation aligns with a key characteristic of GPICL, where learning from context takes precedence over zero-shot generalization. 3. For greater complexity ($n>6$), the ICL horizon clearly extends far beyond $T=4k$, and the improvement in perplexity exhibits polynomial scaling law (depicted as a linear relationship on a semi-logarithmic axis) with increasing context length. However, it remains difficult to assert from current experiments whether, given sufficiently long contexts, models of varying scales will converge to similar asymptotic performances. Nevertheless, we can confidently conclude that for $n<8$, a model with 10 million parameters has been adequate. 4. For the PG-19 test set, we were surprised to find that the tiny-sized model achieves comparable asymptotic performance to the other two models. We must consider that the complexity of natural language far exceeds that of any conceivable $n$-gram language. Furthermore, the improvement does not exhibit a linear trend on a semi-logarithmic scale. It is reasonable to hypothesize that the GPICL is only capturing certain short-term and high-frequency patterns in natural language (such as frequently appeared words).
Our experiments suggest a shift in the GPICL paradigm from “the larger the parameters, the better” to “the longer the context (and the larger the memory), the better”, provided the parameters are sufficient. In contrast, current large language models (LLMs) typically require at least billions of parameters, the majority of which we suspect are employed for zero-shot generalization, while the parameters dedicated to Incremental Context Learning (ICL) are relatively sparse. This perspective is consistent with the concept of the genomics bottleneck (Wang et al., 2022), which advocates for models that, while adequately parameterized, prioritize larger memory capacities to optimize GPICL. However, it is important to note that while the transformer architecture offers a “full memory” capability for sequence modeling, it also incurs a computational cost of $\Theta(T^{2})$. This underscores the necessity for developing more efficient transformers, which would allow for the extension of context length to longer spans. We believe research in this area (Tay et al., 2022) could benefit from this benchmark.
Figure 5: Evaluating language models of various sizes on Meta-Language benchmark with differing levels of complexity ($n$).
3 Learning to navigate an randomized world: Maze World
3.1 Problem Setting
The Maze World simulation tasks an agent with navigating a completely unknown environment, requiring it to explore, memorize its surroundings, perform simultaneous localization and mapping (SLAM), identify navigation targets, and plan its route. This setting could potentially reflect a wide range of real-world tasks, such as household robots adapting to unfamiliar indoor environments. As the agent gains more experience within a specific environment, it should progressively gather more information and navigate to specified targets more efficiently. The proposed benchmarks has the following features:
  • Each task corresponds to a randomly generated $n\times n$ maze, with layouts and textures assigned randomly.
  • A number of Potential Navigation Targets (PNTs) are placed at random locations and remain fixed throughout the task. Each PNT is marked by a semi-transparent beam of light, as shown in Figure 6(a).
  • For each task, a randomized sequence of commands is generated, each directing the agent to walk to a specified target that is identifiable by color.
The agent receives a positive reward for reaching the specified target, while a minor cost is incurred for each step taken. Available actions include moving forward, moving backward, stopping, turning left, and turning right. Beyond these task settings, the benchmarks provide a wide range of hyper-parameters to generate diverse tasks. Examples include the height and angle of the perspective, ceiling height, and the length of the cell unit. In addition to command-based navigation, we also offer a “Survival” mode, details are provided in the appendices.
To optimize the initiation of model training, we have implemented rule-based smart agents, which are allowed to access the global map, but with limitations. They record everything within their line of sight into memories, while areas out of sight remain hidden. This kind of agent efficiently navigates to its goals by utilizing an exploration-then-exploitation strategy. Although this rule-based smart agent may not consistently reach the global optimum, primarily due to shortcomings in its exploration strategy, it nonetheless provides a solid baseline for imitation learning.
To model varied agent behaviors, we differentiate between Long-Term Memory (LTM) and Short-Term Memory (STM) in the rule-based agents. The STM retains only the most recent three steps of observations, while the LTM retains information until the episode concludes. We introduce a hyper-parameter, $p(\text{STM}\rightarrow\text{LTM})$, to govern the probability of transferring information from STM to LTM. At $p(\text{STM}\rightarrow\text{LTM})=0\%$, the agent operates without LTM, whereas $p(\text{STM}\rightarrow\text{LTM})=100\%$ allows for a fully functional LTM.
(a) A demonstration of observations and a global map in MazeWorld
(b) A sketch of an agent’s LTM
Figure 6: A demonstration of the maze world tasks alongside a rule-based navigation system. The rule-based agent is provided with partial access to the environment’s ground truth (b) under specific conditions. Only the areas within the agent’s line of sight are revealed, while the unsighted regions remain obscured. The agent retains the visual information from the nearest three frames in its short-term memory, which is then transferred to long-term memory based on a specified probability, denoted as $p(\textit{STM}\rightarrow\textit{LTM})$ The agent employs a strategy that balances exploration with exploitation.
3.2 Baseline Solution
We offer a baseline solution for the Maze World tasks, which is largely inspired by model-based imitation learning (Hu et al., 2022), causal decision models (Chen et al., 2021), and meta-reinforcement learning (Duan et al., 2016). We use Variational Auto Encoder (VAE) to process each image to latent representation, and use a causal transformer (Chen et al., 2021) to encode the historical observation and action. The model ouputs the distribution of actions, and also the expected next frame. Specifically, given the reference trajectory $(i_{1},a_{1},i_{2},a_{2},...,i_{T},a_{T},i_{T+1})$, where $i_{t}$ and $a_{t}$ represent the observation and action respectively, our model is represented by:
$z_{t}$ $=\textbf{ENC}_{\text{vae}}(i_{t}),$ (2)
$\hat{i}_{t}$ $=\textbf{DEC}_{\text{vae}}(z_{t}),$ (3)
$\pi(a_{0}),\hat{z}_{1},\pi(a_{1}),\hat{z}_{2},...$ $=\textbf{CAUSAL}(z_{0},a_{0},z_{1},a_{1},...),$ (4)
where $z_{t}$ is the latent embedding of the observed image $i_{t}$. Predicting the distribution of actions ($\pi$) and the next frame ($\hat{z}_{t}$) are typically referred as Policy Model and World Model respectively.
We train the model using imitation learning, although reinforcement learning is desirable for future enhancements. The losses are composed of three parts: 1. the reconstruction error ($\mathcal{L}_{vae}$); 2. the cross entropy between the reference action and the predicted distribution ($\mathcal{L}_{pm}$); 3. the next-frame prediction error ($\mathcal{L}_{wm}$). The final loss would be:
$\mathcal{L}_{vae}$ $=\sum_{t}||\hat{i}_{t}-i_{t}||^{2}+\lambda\text{KL}(p(z_{t})||\mathcal{N}(0,1))$ (5)
$\mathcal{L}_{pm}$ $=\sum_{t}-log\pi(a_{t})$ (6)
$\mathcal{L}_{wm}$ $=\sum_{t}||\textbf{DEC}_{\text{vae}}(\hat{z}_{t})-i_{t}||^{2}$ (7)
$\mathcal{L}$ $=\alpha_{1}\mathcal{L}_{vae}+\alpha_{2}\mathcal{L}_{pm}+\alpha_{3}\mathcal{L}_{wm}$ (8)
The choice of causal transformer includes a small-sized transformer (26M parameters) and a standard-sized transformer (330M parameters). A overview of the model architecture is in Figure 7. To train our baseline models, we collect training data include 12.8K episodes of demonstration from rule agents for imitation learning. To avoid the growing “compounding errors” (Ross & Bagnell, 2010) associated with direct behavior cloning, we drew inspiration from data augmentation methods (Ross et al., 2011). We collected the trajectory of states by running a very noisy behavior policy, but used a rule-based agent to label its action at each step (which is not actually executed; see appendices for details). Each episode has $2,048$ steps, which gives $26$M frames in total.
Figure 7: An overview of the model evaluated on the maze world tasks. Initially, a Variational Auto-Encoder (VAE) is trained to encode observed images into hidden vectors and vice versa, with its parameters subsequently frozen. The sequence of observations and actions is then utilized to train a causal model, specifically a transformer that employs only backward attention, using imitation learning and self-supervised learning techniques.
3.3 Baseline Performances
3.4 Static Evaluation
(a) $\mathcal{L}_{wm}(t)$
(b) $\mathcal{L}_{pm}(t)$
Figure 8: Static assessment of the ability to mimic rule-based agents ($\mathcal{L}_{pm}$) and predict the next frame ($\mathcal{L}_{wm}$) across each time step ($t$). Context lengths of 512 and 2048 are evaluated respectively for comparison.
The static evaluation was conducted using off-policy data from an independently sampled validation set comprising $720$ sequences (totaling $1.5$M frames). To analyze the impact of context length on performance, we assessed the losses of the world model and policy model for each individual time step ($\mathcal{L}_{wm}(t)$ and $\mathcal{L}_{pm}(t)$). In Figure 8, We compare the performance of the small-sized and standard-sized causal models, with several noteworthy observations: First, the future prediction task shows no clear dependency on long contexts (the learning horizon is below 100 steps), whereas the performance of action prediction continually improves with longer contexts. This likely suggests that world modeling in the proposed tasks is more related to few-shot ICL capabilities than to GPICL. This is unsurprising, as world modeling is more connected with understanding the localized environment, whereas navigation requires more global reasoning. Second, the standard-sized model does not demonstrate a clear advantage over the small-sized model. This phenomenon aligns with that observed in meta-language tasks, further validating the idea that scaling up parameters is not always necessary compared to scaling up contexts in GPICL.
3.5 Interactive Evaluation and Case Study
(a) Maze $15\times 15$
(b) Maze $25\times 25$
(c) Maze $35\times 35$
Figure 9: Interactive evaluation results of the rule-based agents, randomized policy, and the causal model, depicted by plotting accumulated rewards against simulated time steps. Each plot summarizes the results from identical 32 tasks. The step minor cost is set to 0.
We further investigate the online performance of the model by directly interacting with the environment using the learned model. The evaluation set includes 32 tasks across mazes of sizes $15\times 15$, $25\times 25$, $35\times 35$. A softmax policy, based on the predicted distribution of actions, is implemented. We plot the mean and variance of the average accumulated rewards over time (t). For comparative analysis, we also include the performance of a random policy, where actions are randomly sampled regardless of the observations, and agents with different short-term memory to long-term memory transition probabilities $p(\text{STM}\rightarrow\text{LTM})$ of $5\%$, $25\%$, $100\%$ (Figure 9). The results indicate that our model significantly outperforms random exploration and rule-based agents with low long-term memory capabilities ($p(\text{STM}\rightarrow\text{LTM})=5\%$). However, there is a noticeable gap compared to those with more effective long-term memories ($p(\text{STM}\rightarrow\text{LTM})=25\%,100\%$). Moreover, the rate of improvement is not significant compared to those with effective long-term memories, where we would expect to see an accelerating accumulation of rewards. We believe that there is ample room for improvement by employing more sophisticated models and better learning algorithms, and extending the context length.
Figure 10: Randomly selected clips consisting of four continuous frames from the evaluated tasks (generated by a standard-size causal model) are displayed in sequence from top to bottom. The images from left to right: initial observations, observations reconstructed by the VAE, ground truth observations for the next frame, and predicted next frame.
In Figure 10, we present the performance of next-frame predictions for randomly selected clips of four frames generated by the causal decision model. The causal modeling successfully predicts most of the maze layouts in the next frame (including textures on the walls), but occasionally misses distant details. Those distant details contribute less to losses and are more dependent on long-term observations, making them more challenging to acquire. It is also worth noting that the failure to predict the next command (as seen in the second case) is frequent but expected, since the commands are issued randomly.
4 Related Work
4.1 Meta-learning and GPICL
General purpose in-context learning (GPICL) is closely related to meta-learning , particularly model-based meta-learning (Duan et al., 2016; Santoro et al., 2016). The latter generally encompasses a wide range of methodologies, including those that rely on gradient-based adaptation (Finn et al., 2017), while in-context learning focuses on gradient-free “instruction following” (Ouyang et al., 2022). In recent years, the distinction between meta-learning and pre-training-ICL has become increasingly blurred (Reed et al., 2022; Dai et al., 2023; Li et al., 2023). However, GPICL fundamentally differs from traditional model-based meta-learning and ICL in large-scale language models in the key aspects including minimal zero-shot generalization, long learning horizon, and abundancy of learning algorithm (where supervised learning, reinforcement learning and others are included).
4.2 Benchmarks for Meta-Learning and ICL
A variety of benchmarks have been utilized to evaluate meta-learning algorithms, encompassing domains such as image classification (LeCun et al., 1998), video games, and robot manipulation (Brockman et al., 2016). Many of these benchmarks were not initially designed for meta-learning but have been adapted to accommodate varying tasks through parametric adjustments in hyper-parameters, such as target positions and robot configurations. Meta-World (Yu et al., 2020) introduces a series of robot arm manipulation tasks that incorporate both parametric and non-parametric variations.
A key distinction between our proposed benchmarks and previous ones in meta-learning is that the latter often veer towards "task identification" by selectively omitting critical observations. The setup of these tasks can appear forced and unrealistic. For example, scenarios involve a robot moving in a direction that remains hidden and can only be inferred from rewards (Finn et al., 2017), or a robot arm reaching for a target that is neither visible nor directly observed (Yu et al., 2020). In practical settings, such information like targets and directions would typically be explicitly indicated through instructions or visual cues, making the tasks ostensibly complex yet fundamentally simplistic. Moreover, these tasks generally do not necessitate long-term interaction, with most being solvable through relatively brief explorations. Our benchmarks, in contrast, align more closely with real-world challenges and include tasks such as indoor navigation, planning, Simultaneous Localization and Mapping (SLAM), and language modeling. These are designed to mirror actual scenarios where long-term interaction and complex decision-making are critical.
Another set of benchmarks has emerged alongside large language models (LLMs), focusing on evaluating the capability of modeling long sequences. The Long Range Arena (Tay et al., 2020) includes six tasks, primarily adapting image classification tasks to pixel-wise classification. Other benchmarks involve natural corpora, such as literary texts (Rae et al., 2019) and long document classification (Chalkidis et al., 2019). However, a significant issue with using these benchmarks to study GPICL is that it is often difficult to distinguish between GPICL capabilities and zero-shot generalization. Indeed, previous research on efficient transformers (Tay et al., 2022) has shown that improvements in zero-shot generalization (such as through simply increasing model scale) tend to outperform increases in context length (especially beyond 4,000 tokens) in these datasets. This suggests that these benchmarks may not be ideal for thoroughly testing GPICL.
5 Conclusions
In this paper, we enrich and redefine the topic of General Purpose In-Context Learning (GPICL) and underscore its significance to the AI community. We introduce two benchmarks specifically designed for GPICL. Our preliminary investigations of these benchmarks have yielded some inspiring results, demonstrating that a completely randomized meta-language can be used to train a language model capable of understanding realistic natural language to a certain extent.
It is important to note that although we are seeing the potential of application of those datasets in areas including language modeling and indoor navigation, the synthetic data generated by those benchmarks are not yet validated to be ready for a direct help to application. Currently, large language models (LLMs) typically feature vocabulary sizes $N$ exceeding tens of thousands, highlighting that the true complexity of natural language far surpasses $8$-gram. Moreover, the visual signals in maze worlds remain regularized and relatively monotonic, still significantly less complex than those found in the real world.
This work can be expanded along two directions: 1. The development of more realistic GPICL benchmarks, for instance, by adding greater complexity to the meta-language; 2. Utilizing the benchmarks to uncover more sophisticated models (especially for long-term dependency modeling) and optimization techniques (such as reinforcement learning) to further extend the GPICL horizon and enhance performance gains.

### v2

@@
 Abstract
-In-context learning (ICL) capabilities is becoming increasingly appealing towards building general intelligence.
-Taking this concept one step further, we draw a parallel to humans and many animals, who inherit primarily learning capabilities but refine their memory and acquire diverse skills and knowledge through extensive lifelong experiences.
-This parallel inspires our approach to general purpose in-context learning (GPICL).
-This paper introduces two lightweight but insightful benchmarks specifically crafted to train and evaluate GPICL functionalities.
-Each benchmark encompasses a wide range of diverse tasks characterized by generation and interaction, minimal transferable knowledge, and long-term dependency.
-These features present significant challenges for models that primarily rely on context or interactions to enhance their proficiency.
-We hope that these benchmarks will not only advance research in GPICL but also contribute significantly to the broader field of general intelligence.
+In-context learning (ICL) capabilities are becoming increasingly appealing for building general intelligence due to their sample efficiency and independence from artificial optimization skills.
+To enhance generalization, biological neural systems primarily inherit learning capabilities and subsequently refine their memory, acquiring diverse skills and knowledge through extensive lifelong experiences.
+This process gives rise to the concept of general-purpose in-context learning (GPICL).
+Compared to standard ICL, GPICL addresses a broader range of tasks, extends learning horizons, and starts at a lower zero-shot baseline.
+We introduce two lightweight but insightful benchmarks specifically crafted to train and evaluate GPICL functionalities.
+Each benchmark includes a vast number of tasks characterized by significant task variance and minimal transferable knowledge among tasks, facilitating lifelong in-context learning through continuous generation and interaction.
+These features pose significant challenges for models that rely on context or interactions to improve their proficiency, including language models, decision models, and world models.
+Our experiments reveal that parameter scale alone may not be crucial for ICL or GPICL, suggesting alternative approaches such as increasing the scale of contexts and memory states.
 00footnotetext: Works done at Baidu Inc.
-and University of Science and Technology of China.00footnotetext: Corresponds to: fanwang.px@gmail.com and forrest@ustc.edu.cn00footnotetext: source code available at https://github.com/FutureAGI/L3C
+and University of Science and Technology of China.00footnotetext: Corresponds to: Fan Wang (fanwang.px@gmail.com) and Yang Cao (forrest@ustc.edu.cn)00footnotetext: source code available at https://github.com/FutureAGI/L3C
 1 Introduction
@@
 This disparity is further highlighted by the phenomenon known as the genomics bottleneck.
-Understanding this could guide the design of machine intelligence systems (Zador, 2019; Kirsch & Schmidhuber, 2021; Wang et al., 2022).
+Understanding this could guide the design of machine intelligence systems (Zador, 2019; Kirsch and Schmidhuber, 2021; Wang et al., 2022).
 A promising research area is general purpose in-context learning (GPICL) (Kirsch et al., 2022; Wang et al., 2022).
@@
 However, it raises the question of why we should reduce zero-shot capability, which appears to contradict the prevailing trend in large language models.
-First, inspired by the genomics bottleneck and building on previous work (Kirsch & Schmidhuber, 2021; Wang et al., 2022), one potential strategy to enhance generalizability of the meta-model is to maintain compact meta-parameters that are focused primarily on the functionality of learning itself, rather than to keep overwhelming parameters designed to capture all task-specific knowledge and skills.
+First, inspired by the genomics bottleneck and building on previous work (Kirsch and Schmidhuber, 2021), one potential strategy to enhance generalizability of the meta-model is to maintain compact meta-parameters that are focused primarily on the functionality of learning itself, rather than to keep overwhelming parameters designed to capture all task-specific knowledge and skills.
 Second, the use of abundant synthetic data in pre-training and meta-training, which tends to be non-factual and biased, presents challenges to the usage of the model.
@@
 Isolating the ICL capability alone would be beneficial for inspiring better model architectures and learning algorithms.
-Building on those points, we may futher propose an prospective approach for advancing the development of general intelligence in contemporary settings: Initially, a substantial volume of synthesized data, potentially containing a wealth of nonfactual and biased information, is used exclusively to cultivate the GPICL functionality.
+Building on those points, there can be an prospective approach for advancing the development of general intelligence in contemporary settings: Initially, a substantial volume of synthesized data, potentially containing a wealth of nonfactual and biased information, is used exclusively to cultivate the GPICL functionality.
 Subsequently, a considerable amount of high-quality, unbiased, well-aligned data is employed to enrich the contexts and memory, enabling the agent to achieve capabilities approaching those of humans.
@@
 The primary goal of this paper is neither to introduce a new model or learning methodology, nor to provide an exhaustive survey of the existing ICL systems; rather, we aim to demonstrate through these baselines the potential for these benchmarks to benefit the research of ICL and GPICL.
+One crucial conclusion from these limited experiments may be that parameter scale alone is not crucial for ICL or GPICL.
+This finding aligns with previous research indicating that the size of contexts and memory states can play a more significant role (Kirsch et al., 2022; Wang et al., 2022).
+This insight may highlight alternative approaches beyond merely scaling up models within the current AI community.
 2 Learning to learn an randomized language: Meta-Language
@@
 We analyzed the outputs of the MetaLM across two different context lengths, with the shorter contexts comprising only the first half of the longer contexts.
-The results, depicted in Figure 4, show that MetaLM is capable of learning to memorize and correct both English vocabulary and mathematical equations, validating that its ICL capability is independent of vocabulary settings.
+The results, depicted in Fig.
+4, show that MetaLM is capable of learning to memorize and correct both English vocabulary and mathematical equations, validating that its ICL capability is independent of vocabulary settings.
 Furthermore, the longer the context provided, the fewer errors appeared in the outputs, indicating a clear trend of improvement.
@@
 We have conducted further research into how the scale of parameters affects GPICL capability.
-As illustrated in Figure 5, our findings reveal notable trends: 1.
+As illustrated in Fig.
+5, our findings reveal notable trends: 1.
 Language complexity ($n$) influences the relative performance of different-sized models and affects the dependence on context length.
@@
 The choice of causal transformer includes a small-sized transformer (26M parameters) and a standard-sized transformer (330M parameters).
-A overview of the model architecture is in Figure 7.
+A overview of the model architecture is in Fig.
+7.
 To train our baseline models, we collect training data include 12.8K episodes of demonstration from rule agents for imitation learning.
-To avoid the growing “compounding errors” (Ross & Bagnell, 2010) associated with direct behavior cloning, we drew inspiration from data augmentation methods (Ross et al., 2011).
+To avoid the growing “compounding errors” (Ross and Bagnell, 2010) associated with direct behavior cloning, we drew inspiration from data augmentation methods (Ross et al., 2011).
 We collected the trajectory of states by running a very noisy behavior policy, but used a rule-based agent to label its action at each step (which is not actually executed; see appendices for details).
@@
 3.3 Baseline Performances
-3.4 Static Evaluation
 (a) $\mathcal{L}_{wm}(t)$
@@
 Context lengths of 512 and 2048 are evaluated respectively for comparison.
-The static evaluation was conducted using off-policy data from an independently sampled validation set comprising $720$ sequences (totaling $1.5$M frames).
+Static Evaluation.
+We first conduct static evaluation using off-policy data from an independently sampled validation set comprising $720$ sequences (totaling $1.5$M frames).
 To analyze the impact of context length on performance, we assessed the losses of the world model and policy model for each individual time step ($\mathcal{L}_{wm}(t)$ and $\mathcal{L}_{pm}(t)$).
@@
 This likely suggests that world modeling in the proposed tasks is more related to few-shot ICL capabilities than to GPICL.
-This is unsurprising, as world modeling is more connected with understanding the localized environment, whereas navigation requires more global reasoning.
+This is unsurprising, since single-step world modeling is more connected with understanding the localized environment, whereas navigation requires more global reasoning.
 Second, the standard-sized model does not demonstrate a clear advantage over the small-sized model.
 This phenomenon aligns with that observed in meta-language tasks, further validating the idea that scaling up parameters is not always necessary compared to scaling up contexts in GPICL.
-3.5 Interactive Evaluation and Case Study
 (a) Maze $15\times 15$
@@
 The step minor cost is set to 0.
-We further investigate the online performance of the model by directly interacting with the environment using the learned model.
+Interactive Evaluation. It is more beneficial to investigate the online performance of the policy model and world model in causal modeling by directly interacting with the environment.
+We first directly evaluate the model-free policy by interacting with a set of evaluation tasks.
 The evaluation set includes 32 tasks across mazes of sizes $15\times 15$, $25\times 25$, $35\times 35$.
@@
 We believe that there is ample room for improvement by employing more sophisticated models and better learning algorithms, and extending the context length.
-Figure 10: Randomly selected clips consisting of four continuous frames from the evaluated tasks (generated by a standard-size causal model) are displayed in sequence from top to bottom.
-The images from left to right: initial observations, observations reconstructed by the VAE, ground truth observations for the next frame, and predicted next frame.
-In Figure 10, we present the performance of next-frame predictions for randomly selected clips of four frames generated by the causal decision model.
-The causal modeling successfully predicts most of the maze layouts in the next frame (including textures on the walls), but occasionally misses distant details.
-Those distant details contribute less to losses and are more dependent on long-term observations, making them more challenging to acquire.
-It is also worth noting that the failure to predict the next command (as seen in the second case) is frequent but expected, since the commands are issued randomly.
+Figure 10: Randomly selected clips depicting the future 8 steps using the world model of causal modeling are presented.
+For each clip, the upper section shows the ground truth observations from the current step to 8 steps ahead (from left to right), while the bottom section shows the imagined future up to 8 steps ahead starting from the current observation.
+The actions taken are also labeled with text in the image.
+The future imagination is generally accurate for the first frame but deteriorates over time due to error propagation, illustrating the challenge of causal world modeling.
+We then provide a qualitative evaluation of the performance of world models in Figure 10.
+Unlike $\mathcal{L}_{wm}$, which can only represent the precision of predicting one-step transitions, using the world model to predict the distant future is far more challenging.
+We observe that although our world model can predict the nearest future, its accuracy quickly deteriorates over time.
+This can be attributed to the well-known problem of horizon-square compounding errors.
+Typical improvements include adding noise to the inputs (Hu et al., 2022) and multi-step prediction (Asadi et al., 2019).
+However, compared to many previous problems, Maze World poses a higher challenge to world modeling in two key aspects: First, the variance in observations is more severe (the observation of two consecutive frames can be very different).
+Second, the variance of the tasks requires the world model to perform in-context learning to improve, as each maze is completely different and not transferrable.
 4 Related Work
@@
 Meta-World (Yu et al., 2020) introduces a series of robot arm manipulation tasks that incorporate both parametric and non-parametric variations.
+Mazes are also frequently used to benchmark meta-reinforcement learning (Duan et al., 2016; Mishra et al., 2018), but the scale and diversity of those tasks are not adequate enough to benchmark GPICL.
 A key distinction between our proposed benchmarks and previous ones in meta-learning is that the latter often veer towards "task identification" by selectively omitting critical observations.
-The setup of these tasks can appear forced and unrealistic.
 For example, scenarios involve a robot moving in a direction that remains hidden and can only be inferred from rewards (Finn et al., 2017), or a robot arm reaching for a target that is neither visible nor directly observed (Yu et al., 2020).
@@
 Moreover, these tasks generally do not necessitate long-term interaction, with most being solvable through relatively brief explorations.
-Our benchmarks, in contrast, align more closely with real-world challenges and include tasks such as indoor navigation, planning, Simultaneous Localization and Mapping (SLAM), and language modeling.
-These are designed to mirror actual scenarios where long-term interaction and complex decision-making are critical.
+Our benchmarks, in contrast, align more closely with real-world challenges.
+For instance, the Maze World “navigation” tasks can mirror the adaptive learning process of a household robot adapting to a new indoor scenario, where it needs to simultaneously solve localization, mapping, and planning.
+Another group of Maze World “survival” tasks simulate an agent exploring a new environment to learn about helpful and harmful elements.
 Another set of benchmarks has emerged alongside large language models (LLMs), focusing on evaluating the capability of modeling long sequences.
@@
 The development of more realistic GPICL benchmarks, for instance, by adding greater complexity to the meta-language; 2.
-Utilizing the benchmarks to uncover more sophisticated models (especially for long-term dependency modeling) and optimization techniques (such as reinforcement learning) to further extend the GPICL horizon and enhance performance gains.
+Utilizing the benchmarks to uncover more sophisticated models (including memory-augmented models) and optimization techniques (such as reinforcement learning) to further extend the GPICL horizon and enhance performance gains.

### v5

@@
-Benchmarking General Purpose In-Context Learning
+Benchmarking General-Purpose In-Context Learning
 Abstract
-In-context learning (ICL) capabilities are becoming increasingly appealing for building general intelligence due to their sample efficiency and independence from artificial optimization skills.
-To enhance generalization, biological neural systems primarily inherit learning capabilities and subsequently refine their memory, acquiring diverse skills and knowledge through extensive lifelong experiences.
-This process gives rise to the concept of general-purpose in-context learning (GPICL).
-Compared to standard ICL, GPICL addresses a broader range of tasks, extends learning horizons, and starts at a lower zero-shot baseline.
-We introduce two lightweight but insightful benchmarks specifically crafted to train and evaluate GPICL functionalities.
-Each benchmark includes a vast number of tasks characterized by significant task variance and minimal transferable knowledge among tasks, facilitating lifelong in-context learning through continuous generation and interaction.
-These features pose significant challenges for models that rely on context or interactions to improve their proficiency, including language models, decision models, and world models.
-Our experiments reveal that parameter scale alone may not be crucial for ICL or GPICL, suggesting alternative approaches such as increasing the scale of contexts and memory states.
+In-context learning (ICL) empowers generative models to address new tasks effectively and efficiently on the fly, without relying on any artificially crafted optimization techniques.
+In this paper, we study extending ICL to address a broader range of tasks with an extended learning horizon and higher improvement potential, namely General-Purpose In-Context Learning (GPICL).
+To this end, we introduce two lightweight benchmarks specifically crafted to train and evaluate GPICL functionalities.
+Each benchmark encompasses a vast number of tasks characterized by significant task variance, facilitating meta-training that minimizes inductive bias.
+These tasks are also crafted to promote long-horizon in-context learning through continuous generation and interaction.
+These characteristics necessitate the models to leverage contexts and history interactions to enhance their capabilities, across domains such as language modeling, decision-making, and world modeling.
+Our experiments on the baseline models demonstrate that meta-training with minimal inductive bias and ICL from the ground up is feasible across all the domains we’ve discussed.
+Additionally, our findings indicate that the scale of parameters alone may not be crucial for ICL or GPICL, suggesting alternative approaches such as increasing the scale of contexts and memory states.
 00footnotetext: Works done at Baidu Inc.
-and University of Science and Technology of China.00footnotetext: Corresponds to: Fan Wang (fanwang.px@gmail.com) and Yang Cao (forrest@ustc.edu.cn)00footnotetext: source code available at https://github.com/FutureAGI/L3C
+and University of Science and Technology of China.00footnotetext: Corresponds with: Fan Wang (fanwang.px@gmail.com) and Yang Cao (forrest@ustc.edu.cn)00footnotetext: Source code available at https://github.com/FutureAGI/L3C
 1 Introduction
 The success of large-scale generative language models can be primarily attributed to two key factors: 1.
-Zero-shot generalization derived from the extensive accumulation and storage of knowledge within the model parameters during the pretraining and fine-tuning stages (Kojima et al., 2022); 2.
+Zero-shot generalization derived from the extensive accumulation and storage of knowledge within the model parameters during the pre-training and fine-tuning stages (Kojima et al., 2022); 2.
 In-context learning(ICL) to distill knowledge from the context during the inference stage (Brown et al., 2020).
@@
 GPICL Low High Long
-Figure 1: Conceptual diagram of in-context learning (ICL) and general purpose in-context learning (GPICL)
-The community has long recognized a significant gap between artificial intelligence and biological intelligence, particularly regarding innate abilities and the functionality to continuously learn throughout their lifespans (Parisi et al., 2019).
-For example, the initial limited capabilities of mammal infants stand in stark contrast to their skills in lifelong learning and adaptation.
-This disparity is further highlighted by the phenomenon known as the genomics bottleneck.
-Understanding this could guide the design of machine intelligence systems (Zador, 2019; Kirsch and Schmidhuber, 2021; Wang et al., 2022).
-A promising research area is general purpose in-context learning (GPICL) (Kirsch et al., 2022; Wang et al., 2022).
+Figure 1: Conceptual diagram of in-context learning (ICL) and general-purpose in-context learning (GPICL)
+There are significant gap between artificial intelligence and biological intelligence, particularly regarding the functionality to continuously learn throughout their lifespans (Parisi et al., 2019).
+The enormous lifelong learning and adaptation potential of biological neural networks, forming abrupt contrast with their limited innate abilities at the initial stages (e.g., the stage of mammalian infancy), are also used to guide the design of machine intelligence systems (Zador, 2019; Wang et al., 2022; Schmidgall et al., 2024).
+A promising area in this line of research is general-purpose in-context learning (GPICL) (Kirsch et al., 2022; 2023).
 GPICL involves meta-training across a wide range of task classes and learning to interpret context within a variety of highly diverse tasks.
@@
 Figure 2: Learning pipelines based on GPICL versus the current LLM learning pipeline (Ouyang et al., 2022).
-Why GPICL?
+Importance of GPICL.
 The objective of extending the in-context learning (ICL) horizon and potential is widely acknowledged.
-However, it raises the question of why we should reduce zero-shot capability, which appears to contradict the prevailing trend in large language models.
-First, inspired by the genomics bottleneck and building on previous work (Kirsch and Schmidhuber, 2021), one potential strategy to enhance generalizability of the meta-model is to maintain compact meta-parameters that are focused primarily on the functionality of learning itself, rather than to keep overwhelming parameters designed to capture all task-specific knowledge and skills.
-Second, the use of abundant synthetic data in pre-training and meta-training, which tends to be non-factual and biased, presents challenges to the usage of the model.
-In this context, GPICL could offer an ideal solution: since the meta-training stage can result in the loss of much task-specific knowledge and skills but preserves the core learning functionality, it helps prevent the meta-model from being overly influenced by non-factual information and bias in those tasks.
-Third, in practice, a model’s performance typically depends on both in-context learning (ICL) and zero-shot generalization capabilities, making it challenging to investigate and improve ICL independently.
-Isolating the ICL capability alone would be beneficial for inspiring better model architectures and learning algorithms.
-Building on those points, there can be an prospective approach for advancing the development of general intelligence in contemporary settings: Initially, a substantial volume of synthesized data, potentially containing a wealth of nonfactual and biased information, is used exclusively to cultivate the GPICL functionality.
-Subsequently, a considerable amount of high-quality, unbiased, well-aligned data is employed to enrich the contexts and memory, enabling the agent to achieve capabilities approaching those of humans.
-A visual illustration of this proposed pipeline with existing methodologies is depicted in Figure 2.
+However, suppressing zero-shot capability seem to contradict the prevailing trend in large language models.
+To elucidate its importance further, we emphasize three key points:
+First, theoretically, the lower zero-shot capability in meta-training indicates a reduced reliance on inductive biases, thereby demonstrating sufficient diversity of meta-training tasks.
+As a result, the meta-training process focus primarily on the functionality of ICL instead of task-specific knowledge.
+Conversely, the attainment of zero-shot capability might suggest that the tasks possess a common knowledge base.
+While it could lead to faster adaption within the scope of these tasks, it might potentially restrict the model’s applicability beyond the scope of these tasks.
+Second, practically, meta-training the learning functionality requires an enormous number of tasks, which is challenging to collect in the real world.
+The use of abundant synthetic data in pre-training and meta-training can be essential, but tend to be non-factual and biased, presenting challenges such as the sim-to-real gap.
+Ideally, the meta-training stage of GPICL results in the loss of much task-specific knowledge while preserving the ICL functionality, thus helps prevent the model from the biases in the synthetic datasets.
+Building on this point, an alternative approach could be developed to replace the current mainstream learning pipeline of Pre-training & Fine-tuning & ICL with Meta-training & GPICL (Figure 2).
+In the new pipeline, initially, massive synthesized data, potentially containing a wealth of nonfactual and biased information, is used exclusively to cultivate the GPICL functionality.
+Subsequently, a considerable amount of high-quality, unbiased, well-aligned, real-world data is employed to enrich the contexts and memory, enabling the agent to achieve capabilities approaching those of humans.
+Third, since a model’s performance typically depends on both in-context learning (ICL) and zero-shot generalization capabilities, it is challenging to investigate and improve ICL independently.
+Isolating the ICL capability alone would be beneficial for analysis on model architectures and learning algorithms.
 Guidelines for designing GPICL benchmarks.
@@
 We suggest that those benchmarks should satisfy the following standard:
-  • Vast number of variant tasks with minimal knowledge transfer: To minimize the effectiveness of zero-shot generalization, it is advantageous to engage with a variety of tasks where task-specific knowledge cannot be transferred among different tasks.
-This guarantees that the model’s performance is primarily driven by its in-context learning capabilities rather than its ability of zero-shot generalization.
-It is essential to require numerous tasks to prevent the meta-model from hacking in-context learning (ICL) by merely memorizing all possible tasks and performing task identification, a limitation that most meta-learning benchmarks fall into when given limited tasks.
-  • Interactive and generative tasks: GPICL should include task exploration and completion processes that are both interactive and generative, such that in-context reinforcement learning capabilities can be effectively meta-trained and evaluated.
+  • Vast number of variant tasks with minimal inductive bias: It is desirable to ensure that the meta-training and meta-testing tasks are sufficiently diverse, preventing the transfer of task-specific knowledge among different tasks.
+This guarantees that the model’s performance is primarily driven by its in-context learning capabilities rather than inductive biases.
+It is also essential to require numerous tasks to prevent the model from hacking in-context learning (ICL) by merely memorizing all possible tasks and performing task identification, a limitation that many meta-learning benchmarks fall into when given limited tasks, as is discussed in Kirsch et al.
+(2022).
+  • Interactive and generative tasks: GPICL should include task exploration and completion processes that are both interactive and generative, to ensure obtaining complex functionalities such as in-context reinforcement learning.
 Consequently, benchmarks that solely focus on encoding a single sample characterized by extended contexts and features and conducting single-step predictions or classifications do not meet the criteria.
-  • Lifelong in-context learning: GPICL should handle very long-term dependencies that can scale up to millions or even billions of steps, effectively supporting the learning process over extensive periods (ICL horizon $T$ should be “from cradle to grave”).
-It is necessary to clarify that $T$ should denote the “minimum steps” possible given “ideal ICL functionality”.
-While many previous gradient-based meta-learning methods also encode adaptation over millions of samples (e.g., many gradient-based meta-learning methods), most of these are to blame on the inefficient learning functionality.
-Consequently, they should not be considered as representing a long ICL horizon.
+  • Lifelong in-context learning: To guarantee the capability of adapting complex tasks, GPICL should handle very long-term dependencies that can scale up to millions or even billions of steps, effectively supporting the learning process over extensive periods (ICL horizon $T$ should be “from cradle to grave”).
+It is important to clarify that $T$ should denote the “minimum steps” possible given “ideal ICL functionality”, excluding scenarios requiring extensive contexts that are primarily due to inefficient learning functionalities.
 Contribution of this paper: We propose two critical benchmarks that adhere to the GPICL standard and are capable of generating unlimited synthetic data: 1.
@@
 Maze World: We generate an unlimited number of tasks within randomized mazes, where the model is meta-trained solely through context to learn exploration and navigation.
-For both benchmarks, we provide limited experiment results of baseline models based on Transformer (Vaswani et al., 2017) backbone.
+For both benchmarks, we provide limited experiment results of baseline models based on Transformer (Vaswani et al., 2017) backbone and its variants.
 The primary goal of this paper is neither to introduce a new model or learning methodology, nor to provide an exhaustive survey of the existing ICL systems; rather, we aim to demonstrate through these baselines the potential for these benchmarks to benefit the research of ICL and GPICL.
-One crucial conclusion from these limited experiments may be that parameter scale alone is not crucial for ICL or GPICL.
-This finding aligns with previous research indicating that the size of contexts and memory states can play a more significant role (Kirsch et al., 2022; Wang et al., 2022).
-This insight may highlight alternative approaches beyond merely scaling up models within the current AI community.
+One crucial conclusion from these limited experiments may be that the scale of parameters alone is not crucial for ICL or GPICL.
+This finding aligns with previous research indicating that the size of contexts, and the scale of the memory states can play a more significant role (Kirsch et al., 2022; Wang et al., 2022) in learning from contexts.
 2 Learning to learn an randomized language: Meta-Language
@@
 It is important to note that the model has never been trained on any natural language corpora.
-Specific details are provided in the appendices.
+Details of the model description can be found in Appendix A.1.
 2.3 Evaluation Results
@@
 Most importantly, the model demonstrates a significant performance gain through in-context learning (ICL) in realistic natural language scenarios, despite never being exposed to a single real-world natural word, illustrating that the ICL capability acquired through meta-language is not restricted to a specific language type.
-We test the ability of generation of the meta-language model on two rudimentary real-world language tasks: English vocabulary learning and basic mathematical operations (addition and subtraction), which are typical for primary school education.
-For the English vocabulary task, the vocabulary consists of alphabet characters and common punctuation marks; for the math task, it includes digits and mathematical operators.
-The input contexts for testing included 40 English words and simple arithmetic operations below $5$.
-To discourage direct replication, we intentionally introduced errors into $15\%$ of the words or equations.
-We analyzed the outputs of the MetaLM across two different context lengths, with the shorter contexts comprising only the first half of the longer contexts.
-The results, depicted in Fig.
-4, show that MetaLM is capable of learning to memorize and correct both English vocabulary and mathematical equations, validating that its ICL capability is independent of vocabulary settings.
-Furthermore, the longer the context provided, the fewer errors appeared in the outputs, indicating a clear trend of improvement.
-(a) English vocabulary learning task.
-(b) Mathematical operation task.
-Figure 4: Input contexts were designed such that $15\%$ contained errors.
-All mistakes are highlighted with red squares.
-The vocabulary for both tasks was limited to 32 pre-defined tokens, and there was no pre-training or further tuning applied to either task.
-Notice that this setup tests the MetaLM’s ability to adapt and correct based on in-context learning alone, without relying on prior knowledge or adjustments.
-The model under evaluation has not been trained on any natural langauge corpora.
-Additionally, the vocabulary is randomly assigned to the predefined 32 tokens prior to inference.
+We further test the ability of generation of the meta-language model on two rudimentary real-world language tasks, showing that in contrast to the absence of any natural language corpora, the model is able to learn to adapt to the real world language tasks through ICL only (Appendix A.3).
 2.3.2 Scaling Laws for GPICL
@@
 As illustrated in Fig.
-5, our findings reveal notable trends: 1.
+4, our findings reveal notable trends: 1.
 Language complexity ($n$) influences the relative performance of different-sized models and affects the dependence on context length.
@@
 It is reasonable to hypothesize that the GPICL is only capturing certain short-term and high-frequency patterns in natural language (such as frequently appeared words).
-Our experiments suggest a shift in the GPICL paradigm from “the larger the parameters, the better” to “the longer the context (and the larger the memory), the better”, provided the parameters are sufficient.
+Our experiments suggest a shift in the GPICL paradigm from increasing the parameters to increasing the context length, memory, and states provided the parameters are sufficient.
 In contrast, current large language models (LLMs) typically require at least billions of parameters, the majority of which we suspect are employed for zero-shot generalization, while the parameters dedicated to Incremental Context Learning (ICL) are relatively sparse.
@@
 3.1 Problem Setting
-The Maze World simulation tasks an agent with navigating a completely unknown environment, requiring it to explore, memorize its surroundings, perform simultaneous localization and mapping (SLAM), identify navigation targets, and plan its route.
+Mazes are typical environments that can scale up to an unlimited number of layouts, features, targets, and tasks.
+Many previous works benchmark meta-learning with mazes, but these are limited to 2-D observation space (Wang et al., 2022; Morad et al., 2023; Nikulin et al., 2023), text-only (Ding et al., 2024) observation and action space, and limited number of layouts and tasks (Duan et al., 2016; Morad et al., 2023; Ding et al., 2024).
+To enhance the task diversity, we propose the Maze World benchmark.
+It tasks an agent with navigating a completely unknown environment, requiring it to explore, memorize its surroundings, perform simultaneous localization and mapping (SLAM), identify navigation targets, and plan its route.
 This setting could potentially reflect a wide range of real-world tasks, such as household robots adapting to unfamiliar indoor environments.
@@
   • For each task, a randomized sequence of commands is generated, each directing the agent to walk to a specified target that is identifiable by color.
-The agent receives a positive reward for reaching the specified target, while a minor cost is incurred for each step taken.
+The agent receives a positive reward for reaching the specified target, while a minor step-cost is incurred for each step taken.
 Available actions include moving forward, moving backward, stopping, turning left, and turning right.
-Beyond these task settings, the benchmarks provide a wide range of hyper-parameters to generate diverse tasks.
-Examples include the height and angle of the perspective, ceiling height, and the length of the cell unit.
-In addition to command-based navigation, we also offer a “Survival” mode, details are provided in the appendices.
-To optimize the initiation of model training, we have implemented rule-based smart agents, which are allowed to access the global map, but with limitations.
-They record everything within their line of sight into memories, while areas out of sight remain hidden.
-This kind of agent efficiently navigates to its goals by utilizing an exploration-then-exploitation strategy.
-Although this rule-based smart agent may not consistently reach the global optimum, primarily due to shortcomings in its exploration strategy, it nonetheless provides a solid baseline for imitation learning.
-To model varied agent behaviors, we differentiate between Long-Term Memory (LTM) and Short-Term Memory (STM) in the rule-based agents.
-The STM retains only the most recent three steps of observations, while the LTM retains information until the episode concludes.
+Note that unlike previous works (Chen et al., 2021; Kirsch et al., 2023; Lin et al., 2023; Lee et al., 2024), we did not use rewards as inputs, nor did we explicitly predict rewards.
+In the “NAVIGATION” tasks of Maze World, instant rewards are easily inferred from observations.
+However, in other tasks within Maze World, such as “SURVIVAL” tasks (see details in the Appendix A.5), modeling rewards may be necessary.
+This should not impede the analysis of the benchmark and GPICL process.
+To provide a cost-effective and scalable reference dataset for meta-training GPICL, we have implemented rule-based agents with a robust simulated policy.
+These agents are allowed to access the global maps, but with limitations: They are allowed to record map areas within their line of sight into memory, while other areas remain hidden.
+These agents are referred to as privileged agents because they have access to additional information.
+The privileged agents efficiently navigates to their goals by utilizing an exploration-then-exploitation strategy.
+Although these agents may not consistently reach the global optimum due to shortcomings in its exploration strategy, it provides a relatively high standard reference policy for imitation learning.
+To simulate variant levels of intelligent agents, we equip the privileged agents with both Long-Term Memory (LTM) and Short-Term Memory (STM).
+The STM retains only the most recent three pieces of observation, while the LTM retains information until the episode concludes.
+Only part of the STM is allowed to be recorded to LTM.
 We introduce a hyper-parameter, $p(\text{STM}\rightarrow\text{LTM})$, to govern the probability of transferring information from STM to LTM.
 At $p(\text{STM}\rightarrow\text{LTM})=0\%$, the agent operates without LTM, whereas $p(\text{STM}\rightarrow\text{LTM})=100\%$ allows for a fully functional LTM.
+For simplicity, a privileged agent with $\text{p}$ ( STM → LTM ) = p is written as preiviledged agent (p) for short.
 (a) A demonstration of observations and a global map in MazeWorld
-(b) A sketch of an agent’s LTM
-Figure 6: A demonstration of the maze world tasks alongside a rule-based navigation system.
-The rule-based agent is provided with partial access to the environment’s ground truth (b) under specific conditions.
+(b) The LTM of a privileged agent
+Figure 5: A demonstration of the maze world tasks alongside privileged agents with smart navigation policy.
+The privileged agent is provided with limited access to the environment’s ground truth (b) under specific conditions.
 Only the areas within the agent’s line of sight are revealed, while the unsighted regions remain obscured.
-The agent retains the visual information from the nearest three frames in its short-term memory, which is then transferred to long-term memory based on a specified probability, denoted as $p(\textit{STM}\rightarrow\textit{LTM})$ The agent employs a strategy that balances exploration with exploitation.
+The agent retains the visual information from the nearest three frames in its short-term memory, which is then transferred to long-term memory based on a specified probability.
+The privileged agent employs an exploration-then-exploitation strategy.
 3.2 Baseline Solution
-We offer a baseline solution for the Maze World tasks, which is largely inspired by model-based imitation learning (Hu et al., 2022), causal decision models (Chen et al., 2021), and meta-reinforcement learning (Duan et al., 2016).
+The baseline solution for the Maze World tasks is inspired by model-based imitation learning (Hu et al., 2022) and causal decision models (Chen et al., 2021).
+Although running meta-reinforcement learning (MetaRL) could be beneficial, we did not directly apply RL due to the significant challenges in scaling RL to meet the required needs.
+Following previous work that suggests imitating an expert as adequate to generate reinforcement learning capabilities (Lin et al., 2023; Zisman et al., 2023; Lee et al., 2024), we mainly focus on imitation learning augmented with world models.
 We use Variational Auto Encoder (VAE) to process each image to latent representation, and use a causal transformer (Chen et al., 2021) to encode the historical observation and action.
@@
 Predicting the distribution of actions ($\pi$) and the next frame ($\hat{z}_{t}$) are typically referred as Policy Model and World Model respectively.
-We train the model using imitation learning, although reinforcement learning is desirable for future enhancements.
-The losses are composed of three parts: 1.
+The learning losses are composed of three parts: 1.
 the reconstruction error ($\mathcal{L}_{vae}$); 2.
@@
 $\mathcal{L}_{wm}$ $=\sum_{t}||\textbf{DEC}_{\text{vae}}(\hat{z}_{t})-i_{t}||^{2}$ (7)
-$\mathcal{L}$ $=\alpha_{1}\mathcal{L}_{vae}+\alpha_{2}\mathcal{L}_{pm}+\alpha_{3}\mathcal{L}_{wm}$ (8)
-The choice of causal transformer includes a small-sized transformer (26M parameters) and a standard-sized transformer (330M parameters).
+$\mathcal{L}$ $=\alpha_{vae}\mathcal{L}_{vae}+\alpha_{pm}\mathcal{L}_{pm}+\alpha_{wm}\mathcal{L}_{wm}$ (8)
+The choice of causal transformer includes a small-sized transformer (26M parameters) and a standard-sized transformer (237M parameters).
 A overview of the model architecture is in Fig.
-7.
-To train our baseline models, we collect training data include 12.8K episodes of demonstration from rule agents for imitation learning.
-To avoid the growing “compounding errors” (Ross and Bagnell, 2010) associated with direct behavior cloning, we drew inspiration from data augmentation methods (Ross et al., 2011).
-We collected the trajectory of states by running a very noisy behavior policy, but used a rule-based agent to label its action at each step (which is not actually executed; see appendices for details).
-Each episode has $2,048$ steps, which gives $26$M frames in total.
+6.
+To train our baseline models, we collect training data include $120$K episodes of demonstration from demonstration of privileged agents for imitation learning.
+Each episode has $2,048$ steps, which gives $240$M frames in total.
+To avoid the growing “compounding errors” (Ross and Bagnell, 2010) associated with direct behavior cloning, we drew inspiration from dataset aggregation (DAgger) methods (Ross et al., 2011).
+We collected the trajectory of states by running a very noisy behavior policy, but used the privileged agent (100%) to label its action at each step (which is not actually executed).
+Additionally, to enhance the stability of auto-regression, we dropout and randomize $15\%$ of the input observations to the causal transformer during training (Details in Appendix A.2).
 Figure 6: An overview of the model evaluated on the maze world tasks.
@@
 The sequence of observations and actions is then utilized to train a causal model, specifically a transformer that employs only backward attention, using imitation learning and self-supervised learning techniques.
+For context-free and partial-context solutions, we intentionally mask some of the backward connections.
+To investigate the impact of context length on performance, we introduce additional two variants that share the same structure as the small-sized causal transformer (with $26$M parameters), but have limited access to the contexts.
+The partial-context causal transformer has an attention window of 2, providing effective context length up to 9.
+The context-free transformer has an attention window of 1, providing effective context length of 1.
+The context-free and partial-context transformers are trained by further fine-tuning on the foundation of the full-context version (Appendix A.1).
 3.3 Baseline Performances
@@
 (b) $\mathcal{L}_{pm}(t)$
-Figure 8: Static assessment of the ability to mimic rule-based agents ($\mathcal{L}_{pm}$) and predict the next frame ($\mathcal{L}_{wm}$) across each time step ($t$).
-Context lengths of 512 and 2048 are evaluated respectively for comparison.
+Figure 7: Static assessment of the ability to imitate previleged agents ($\mathcal{L}_{pm}$) and predict the next frame ($\mathcal{L}_{wm}$) across each time step ($t$).
 Static Evaluation.
-We first conduct static evaluation using off-policy data from an independently sampled validation set comprising $720$ sequences (totaling $1.5$M frames).
+We first conduct static evaluation using off-policy data from an independently sampled validation set comprising $1275$ sequences (totaling $2.6$M frames).
 To analyze the impact of context length on performance, we assessed the losses of the world model and policy model for each individual time step ($\mathcal{L}_{wm}(t)$ and $\mathcal{L}_{pm}(t)$).
-In Figure 8, We compare the performance of the small-sized and standard-sized causal models, with several noteworthy observations: First, the future prediction task shows no clear dependency on long contexts (the learning horizon is below 100 steps), whereas the performance of action prediction continually improves with longer contexts.
-This likely suggests that world modeling in the proposed tasks is more related to few-shot ICL capabilities than to GPICL.
-This is unsurprising, since single-step world modeling is more connected with understanding the localized environment, whereas navigation requires more global reasoning.
-Second, the standard-sized model does not demonstrate a clear advantage over the small-sized model.
-This phenomenon aligns with that observed in meta-language tasks, further validating the idea that scaling up parameters is not always necessary compared to scaling up contexts in GPICL.
+In Figure 7, We compare the performance of the four variants, with several noteworthy observations: First, the ability to encode long contexts significantly influences performance, while the scale of the parameters shows no observable effect, which is also consistent with the conclusion of Meta-Language.
+Second, both world-modeling and policy-modeling improve with increased context length, though they exhibit some subtle differences.
+The capability of world-modeling improves rapidly within the first 20 steps, after which the rate of improvement slows.
+This could be because single-step world modeling is primarily impacted by the localized environment, explaining the rapid initial improvement with short contexts, but is less influenced by distant layouts, leading to slower improvements over longer contexts.
+In contrast, the improvement in policy modeling with increased context length is not monotonic, with the loss remaining relatively unchanged between $10$ and $500$ steps.
+The decision modeling loss is closely related to the reference policy, which follows an exploration-then-exploitation strategy.
+These results may indicate that imitating the exploration strategy is more challenging.
 (a) Maze $15\times 15$
@@
 (c) Maze $35\times 35$
-Figure 9: Interactive evaluation results of the rule-based agents, randomized policy, and the causal model, depicted by plotting accumulated rewards against simulated time steps.
-Each plot summarizes the results from identical 32 tasks.
+Figure 8: Interactive evaluation results of the randomized, previleged and casual-modeling based agents, depicted by plotting accumulated rewards against simulated time steps.
 The step minor cost is set to 0.
-Interactive Evaluation. It is more beneficial to investigate the online performance of the policy model and world model in causal modeling by directly interacting with the environment.
-We first directly evaluate the model-free policy by interacting with a set of evaluation tasks.
-The evaluation set includes 32 tasks across mazes of sizes $15\times 15$, $25\times 25$, $35\times 35$.
-A softmax policy, based on the predicted distribution of actions, is implemented.
-We plot the mean and variance of the average accumulated rewards over time (t).
-For comparative analysis, we also include the performance of a random policy, where actions are randomly sampled regardless of the observations, and agents with different short-term memory to long-term memory transition probabilities $p(\text{STM}\rightarrow\text{LTM})$ of $5\%$, $25\%$, $100\%$ (Figure 9).
-The results indicate that our model significantly outperforms random exploration and rule-based agents with low long-term memory capabilities ($p(\text{STM}\rightarrow\text{LTM})=5\%$).
-However, there is a noticeable gap compared to those with more effective long-term memories ($p(\text{STM}\rightarrow\text{LTM})=25\%,100\%$).
-Moreover, the rate of improvement is not significant compared to those with effective long-term memories, where we would expect to see an accelerating accumulation of rewards.
-We believe that there is ample room for improvement by employing more sophisticated models and better learning algorithms, and extending the context length.
-Figure 10: Randomly selected clips depicting the future 8 steps using the world model of causal modeling are presented.
-For each clip, the upper section shows the ground truth observations from the current step to 8 steps ahead (from left to right), while the bottom section shows the imagined future up to 8 steps ahead starting from the current observation.
-The actions taken are also labeled with text in the image.
-The future imagination is generally accurate for the first frame but deteriorates over time due to error propagation, illustrating the challenge of causal world modeling.
-We then provide a qualitative evaluation of the performance of world models in Figure 10.
-Unlike $\mathcal{L}_{wm}$, which can only represent the precision of predicting one-step transitions, using the world model to predict the distant future is far more challenging.
-We observe that although our world model can predict the nearest future, its accuracy quickly deteriorates over time.
-This can be attributed to the well-known problem of horizon-square compounding errors.
-Typical improvements include adding noise to the inputs (Hu et al., 2022) and multi-step prediction (Asadi et al., 2019).
-However, compared to many previous problems, Maze World poses a higher challenge to world modeling in two key aspects: First, the variance in observations is more severe (the observation of two consecutive frames can be very different).
-Second, the variance of the tasks requires the world model to perform in-context learning to improve, as each maze is completely different and not transferrable.
+Each plot summarizes the results from static 64 tasks.
+The shaded area represents the 95% confidence interval of the expectation.
+Interactive Evaluation.
+It is more beneficial to investigate the online performance of the policy model and world model by directly interacting with the environment.
+We first directly evaluate the performance of the policy model by interacting with a set of evaluation tasks.
+The evaluation set includes $64$ tasks across mazes of sizes $15\times 15$, $25\times 25$, $35\times 35$.
+We plot the average accumulated rewards and their confidence over time (t).
+For comparative analysis, we also include the performance of a random policy, where actions are randomly sampled regardless of the observations, as well as privileged agents with $p(\text{STM}\rightarrow\text{LTM})$ of $5\%$, $25\%$, $100\%$.
+The results (Figure 8) indicate that the model with full contexts outperforms the random policy, context-free agents, and privileged agents ($5\%$), but there is a significant gap compared to privileged agents with $25\%$ and $100\%$ LTM.
+However, when compared with partial-context causal transformers, the superiority of the full-context model is not obvious.
+Considering the conclusion in static evaluation, it is reasonable to consider that the performance of interactive evaluation is constrained by imitation learning itself.
+We believe it is beneficial to incorporate on-policy trajectories in the dataset in the future, either in reinforcement learning or DAgger (Ross et al., 2011).
+(a) Maze $15\times 15$
+(b) Maze $25\times 25$
+(c) Maze $35\times 35$
+Figure 9: Results regarding the in-context improvement of world modeling in Maze World.
+We examine the compound error associated with a 9-step auto-regressive future predicting process, using the learned world models that differ in the context length ($t\in\{1,100,1000,2000\}$).
+The figures presented compare different models to the context-free world model as the baseline.
+The horizontal axis measures the number of imagined steps, with each curve illustrating a distinct context length.
+The mean square error is averaged across 64 tasks, and the shaded area provides a 95% confidence interval for these measurements.
+In-Context Improvement of World Modeling.
+Figure 9 explores the performance of world models in interactive evaluations.
+At steps $1$, $100$, $1000$, and $2000$ of each task, the world model is tasked with “imagining” $9$ steps into the future.
+For the smaller $15\times 15$ mazes, we observe a significant improvement in world modeling as context length increases from $1$ to $1000$, evidenced by a continuous decrease in prediction error.
+However, this improvement plateaus between $1000$ and $2000$ steps.
+With sufficient context, the cumulative error over 9 steps remains lower than that of an average guess, which is the performance of the context-free model.
+For larger mazes of $25\times 25$ and beyond, and when predicting more than 5 steps into the future, none of the trained models outperform an average guess.
+Additionally, the partial-context model does not enhance its predictive capabilities as context length increases, which aligns with expectations.
+The results indicate that the proposed model demonstrates a reasonable capacity to utilize contextual information for future predictions, yet there is considerable room for improvement, especially in solving larger-scale mazes and the efficient utilization of long-term dependencies.
+For further insights and cases, refer to Appendix A.4.
 4 Related Work
 4.1 Meta-learning and GPICL
-General purpose in-context learning (GPICL) is closely related to meta-learning , particularly model-based meta-learning (Duan et al., 2016; Santoro et al., 2016).
-The latter generally encompasses a wide range of methodologies, including those that rely on gradient-based adaptation (Finn et al., 2017), while in-context learning focuses on gradient-free “instruction following” (Ouyang et al., 2022).
-In recent years, the distinction between meta-learning and pre-training-ICL has become increasingly blurred (Reed et al., 2022; Dai et al., 2023; Li et al., 2023).
-However, GPICL fundamentally differs from traditional model-based meta-learning and ICL in large-scale language models in the key aspects including minimal zero-shot generalization, long learning horizon, and abundancy of learning algorithm (where supervised learning, reinforcement learning and others are included).
-4.2 Benchmarks for Meta-Learning and ICL
+Meta-learning encompasses a wide range of methodologies, including gradient-based adaptation (Finn et al., 2017), attention-based adaptation (Mishra et al., 2018), and memory-based adaptation (Santoro et al., 2016; Duan et al., 2016; Wang et al., 2022; Lu et al., 2024).
+Both attention-based and memory-based adaptations are closely related to in-context learning, especially general-purpose ICL.
+Since the discovery that large-scale language models (LLMs) can perform few-shot learning (Ouyang et al., 2022), Transformer architectures and LLMs are increasingly becoming the state-of-the-art for meta-learning and in-context learning (ICL).
+This includes applications ranging from few-shot supervised learning (Min et al., 2021; Garg et al., 2022; Dai et al., 2023; Li et al., 2023), in-context reinforcement learning (Chen et al., 2021; Laskin et al., 2022; Lin et al., 2023), to multi-modal reasoning and embodied control (Reed et al., 2022; Szot et al., 2023).
+However, typical pre-training models only address relatively simple few-shot learning tasks, while more complex tasks still require task-specific or class-specific training.
+It is desirable to introduce general-purpose in-context learning (Kirsch et al., 2022; 2023) for generalizing across a wider variety of tasks, which necessitates benchmarks that can scale up in both quantity and diversity.
+4.2 Relative Benchmarks
+Meta-Learning Benchmarks.
 A variety of benchmarks have been utilized to evaluate meta-learning algorithms, encompassing domains such as image classification (LeCun et al., 1998), video games, and robot manipulation (Brockman et al., 2016).
-Many of these benchmarks were not initially designed for meta-learning but have been adapted to accommodate varying tasks through parametric adjustments in hyper-parameters, such as target positions and robot configurations.
-Meta-World (Yu et al., 2020) introduces a series of robot arm manipulation tasks that incorporate both parametric and non-parametric variations.
-Mazes are also frequently used to benchmark meta-reinforcement learning (Duan et al., 2016; Mishra et al., 2018), but the scale and diversity of those tasks are not adequate enough to benchmark GPICL.
-A key distinction between our proposed benchmarks and previous ones in meta-learning is that the latter often veer towards "task identification" by selectively omitting critical observations.
-For example, scenarios involve a robot moving in a direction that remains hidden and can only be inferred from rewards (Finn et al., 2017), or a robot arm reaching for a target that is neither visible nor directly observed (Yu et al., 2020).
-In practical settings, such information like targets and directions would typically be explicitly indicated through instructions or visual cues, making the tasks ostensibly complex yet fundamentally simplistic.
-Moreover, these tasks generally do not necessitate long-term interaction, with most being solvable through relatively brief explorations.
-Our benchmarks, in contrast, align more closely with real-world challenges.
-For instance, the Maze World “navigation” tasks can mirror the adaptive learning process of a household robot adapting to a new indoor scenario, where it needs to simultaneously solve localization, mapping, and planning.
-Another group of Maze World “survival” tasks simulate an agent exploring a new environment to learn about helpful and harmful elements.
-Another set of benchmarks has emerged alongside large language models (LLMs), focusing on evaluating the capability of modeling long sequences.
-The Long Range Arena (Tay et al., 2020) includes six tasks, primarily adapting image classification tasks to pixel-wise classification.
-Other benchmarks involve natural corpora, such as literary texts (Rae et al., 2019) and long document classification (Chalkidis et al., 2019).
-However, a significant issue with using these benchmarks to study GPICL is that it is often difficult to distinguish between GPICL capabilities and zero-shot generalization.
-Indeed, previous research on efficient transformers (Tay et al., 2022) has shown that improvements in zero-shot generalization (such as through simply increasing model scale) tend to outperform increases in context length (especially beyond 4,000 tokens) in these datasets.
-This suggests that these benchmarks may not be ideal for thoroughly testing GPICL.
+Many of these benchmarks were not initially designed for meta-learning but have been adapted to accommodate varying tasks through parametric adjustments in hyperparameters and by artificially hiding certain parameters from observation.
+This can be further advanced to include precedurely generated random targets of control (Finn et al., 2017; Mishra et al., 2018; Yu et al., 2020; Morad et al., 2023), video games of different difficulty level (Cobbe et al., 2019), geometry of locomotion (Najarro and Risi, 2020), rules of transition (Nikulin et al., 2023), and layouts of mazes (Mishra et al., 2018; Morad et al., 2023).
+However, some of these hidden configurations are relatively simple and fail to generate sufficient diversity, causing models to lean towards acquiring zero-shot and task identification abilities rather than in-context learning from scratch.
+For example, while many previous benchmarks include hiding the target of locomotion, an ideal learning algorithm should be able to locate the hidden target in at most three steps by comparing the rewards of walking in different directions, given that the location of the target is only in two dimensions.
+Another promising approach is to randomize labels (Kirsch et al., 2022; Wei et al., 2023), observations (Morad et al., 2023), and actions (Sinii et al., 2023), which tend to have higher dimensions.
+Based on the prior studies, it is plausible to infer that the diversity of tasks and the resulting generalization capabilities are related to the number of hyperparameters that can be varied.
+In contrast to the artificial randomization of rewards, observations, and targets, we believe that Meta-Language and MazeWorld offer environments that are inherently more complex and significant for benchmarking GPICL.
+Language Modeling.
+Recently, emphasis has been placed on long-term dependency modeling in large language models (LLMs).
+Most of the newly proposed benchmarks support context lengths over 8K tokens (Rae et al., 2019; Tay et al., 2020; Bai et al., 2023; Shen et al., 2023; Song et al., 2024).
+However, there are several limitations when using current language-based benchmarks for in-context learning (ICL) investigation.
+First, some benchmarks focus on one-step classification with very long feature descriptions (Tay et al., 2020; Chalkidis et al., 2019; Bai et al., 2023), which is not suitable for examining the interactive performance of the model.
+Second, it is unclear how much long-term dependency actually exists in these long contexts (Chen et al., 2024).
+Even if some long-term dependencies are present, they may be weak and relatively simplistic (such as retrieval and copy tasks), and could be overshadowed by the powerful zero-shot generalization capabilities of the models.
+This overshadowing can obscure crucial information important for GPICL.
+In practice, it is often found that scaling up model parameters to enhance zero-shot generalization typically yields better results than extending the ICL horizon.
+Visual Navigation.
+There are a few datasets featuring indoor and outdoor robot-environment interaction scenes, reconstructed from images taken in real-world settings (Song et al., 2017; Chang et al., 2017; Xia et al., 2018), or simulated environments (Szot et al., 2021).
+Currently, most of these environments primarily facilitate the learning of zero-shot generalization.
+A significant limitation of these datasets is the high data collection cost and the challenge of scaling up.
+Compared to these benchmarks, the proposed Maze World has a much larger sim-to-real gap, but offers greater diversity and complexity, and is much easier to scale up.
+Additionally, we suggest that reality-generated data could better serve as contexts in the GPICL stage, as illustrated in Figure 2(b), rather than during meta-training.
 5 Conclusions
-In this paper, we enrich and redefine the topic of General Purpose In-Context Learning (GPICL) and underscore its significance to the AI community.
-We introduce two benchmarks specifically designed for GPICL.
-Our preliminary investigations of these benchmarks have yielded some inspiring results, demonstrating that a completely randomized meta-language can be used to train a language model capable of understanding realistic natural language to a certain extent.
+In this paper, we introduce two benchmarks specifically designed for GPICL.
+Our preliminary investigations of these benchmarks have yielded inspiring results, demonstrating that scaling laws might exist between context length and performances, and that the performance is not entirely dependent on the scaling of parameters.
 It is important to note that although we are seeing the potential of application of those datasets in areas including language modeling and indoor navigation, the synthetic data generated by those benchmarks are not yet validated to be ready for a direct help to application.
-Currently, large language models (LLMs) typically feature vocabulary sizes $N$ exceeding tens of thousands, highlighting that the true complexity of natural language far surpasses $8$-gram.
-Moreover, the visual signals in maze worlds remain regularized and relatively monotonic, still significantly less complex than those found in the real world.
 This work can be expanded along two directions: 1.
@@
 Utilizing the benchmarks to uncover more sophisticated models (including memory-augmented models) and optimization techniques (such as reinforcement learning) to further extend the GPICL horizon and enhance performance gains.
+Appendix A Appendix
+A.1 Model Description
+Table 1: Comprehensive configurations of the baseline models
+Model parameters layers $d_{model}$ $n_{heads}$ batch size (tokens)
+Meta-Language Model (tiny) $303$K 6 64 4 $32$K
+Meta-Language Model (small) $9.5$M 12 256 8 $32$K
+Meta-Langauge Model (standard) $151$M 12 1,024 16 $64$K
+Causal Transformer (small) $26$M 8 512 8 $32$K
+Causal Transformer (standard) $237$M 12 1,280 16 $32$K
+Figure 10: Structures of the encoder and decoder in the Variational Autoencoder
+The comprehensive configurations of the baseline models are outlined in Table 1.
+Notably, within the Maze World framework, each temporal increment is encoded by a pair of tokens: an observation and its corresponding action.
+For the optimization of the Meta-Language and Maze World baseline models, we employed the Noam decay scheduler [Vaswani et al., 2017].
+This scheduler peaks at a learning rate of $10^{-3}$ following an initial warm-up phase spanning $1,000$ steps.
+For the image encoder (11M parameters) and image decoder (14M parameters), we employ convolutional and deconvolutional layers augmented with ResNet blocks [He et al., 2016].
+The detailed structure of the VAE is depicted in Figure 10.
+Initially, we train the VAE in isolation by setting $\alpha_{wm}$ and $\alpha_{pm}$ in Equation 8 to 0.
+Subsequently, we freeze the parameters of the encoder and decoder and train the causal transformer only.
+For partial-context and context-free causal transformers, we modify the attention window to 2 and 1, respectively, as shown in Figure 11.
+Since transformers can propagate information between layers, an attention window of 2 results in an effective context size of the number of layers plus 1.
+(a) Full-context
+(b) Partial-context
+(c) Context-free
+Figure 11: Variant attention masks of causal modeling
+(a) Meta-Language Model
+(b) World Modeling
+(c) Policy Modeling
+(d) World Modeling
+(e) Policy Modeling
+Figure 12: The training loss with respect to the steps of optimization. Notice that the partial-context and context-free causal transformers are initialized from the trained full-context transformers.
+Figure 12 presents the training loss versus the number of iterations for the Meta-Language and Maze World baseline models.
+During training baselines for Meta-Language tasks, we observed that standard-sized models converge relatively slowly on the training set.
+To address this, we conducted additional warm-up epochs on a simpler pre-training dataset with $n=2,3$ before transitioning to the main pre-training dataset.
+During training the baselines for Maze World tasks, we found there was mutual interference between the world modeling loss ($\mathcal{L}_{wm}$) and policy modeling loss ($\mathcal{L}_{pm}$), especially in the initial stages.
+Our meta-training underwent several phases where we continuously adjusted the loss weights $\alpha_{wm}$ and $\alpha_{pm}$ in Equation 8.
+To further accelerate and stabilize the convergence of the world model, in addition to these losses, we introduced an auxiliary loss $\mathcal{L}_{z}=\sum_{t}||\hat{z}_{t}-z_{t}||^{2}$ with a weight of $\alpha_{z}=0.05$.
+During the final stage, the hyperparameters are adjusted to $\alpha_{pm}=0.25$ and $\alpha_{wm}=0.70$.
+However, we recommend starting with $\alpha_{wm}>0.90$ and $\alpha_{pm}<0.10$ as the policy model training loss is more volatile initially.
+The context-free and partial-context causal transformers are initialized from the full-context causal transformer, and the hyperparameters remain unchanged.
+A.2 Data Collection Strategy for Maze World
+For dataset collection and static evaluation of Maze World, inspired by dataset aggregation [Ross et al., 2011] and noise distillation [Zisman et al., 2023], we utilize distinct behavior and reference policies.
+The behavior policy is determined by previleged agent where $p(\text{STM}\rightarrow\text{LTM})$ randomly sampled between $[0,50\%]$.
+Additionally, with the probability of $\epsilon\in[0,80\%]$ the behavior action is determined by random decision.
+With the states generated by behavior policy, we use the reference policy with $p(\text{STM}\rightarrow\text{LTM})=100\%$ to label the action.
+We’ve also randomly dropped 20% of the input observations ($z_{t}$) to enhance the robustness of auto-regression.
+Among these, 10% of the input observations are replaced with a special mask token with trainable embeddings, while the remaining 10% are perturbed with Gaussian noise $\mathcal{N}(0,1)$.
+A.3 Additional Results in Meta-Language Model
+(a) English vocabulary learning task.
+(b) Mathematical operation task.
+Figure 13: Input contexts were designed such that $15\%$ contained errors.
+All mistakes are highlighted with red squares.
+The vocabulary for both tasks was limited to 32 pre-defined tokens, and there was no pre-training or further tuning applied to either task.
+Notice that this setup tests the MetaLM’s ability to adapt and correct based on in-context learning alone, without relying on prior knowledge or adjustments.
+The model under evaluation has not been trained on any natural langauge corpora.
+Additionally, the vocabulary is randomly assigned to the predefined 32 tokens prior to inference.
+We test the ability of generation of the meta-language model on two rudimentary real-world language tasks: English vocabulary learning and basic mathematical operations (addition and subtraction), which are typical for primary school education.
+For the English vocabulary task, the vocabulary consists of alphabet characters and common punctuation marks; for the math task, it includes digits and mathematical operators.
+The input contexts for testing included 40 English words and simple arithmetic operations below $5$.
+To discourage direct replication, we intentionally introduced errors into $15\%$ of the words or equations.
+We analyzed the outputs of the MetaLM across two different context lengths, with the shorter contexts comprising only the first half of the longer contexts.
+The results, depicted in Fig.
+13, show that MetaLM is capable of learning to memorize and correct both English vocabulary and mathematical equations, validating that its ICL capability is independent of vocabulary settings.
+Furthermore, the longer the context provided, the fewer errors appeared in the outputs, indicating a clear trend of improvement.
+A.4 Additional Results in Maze World Tasks
+(a) Random
+(b) Context-free
+(c) Partial-context
+(d) Full-context
+(e) Privileged Agent
+Figure 14: Performances of different methods in 3 randomly picked $35\times 35$ tasks.
+Green trajectories indicate time steps $t$ close to 0, while red trajectories represent $t$ close to $T$.
+From left to right, the columns are generated by: (a) Random Policy (b) Context-free Model (c) Partial-context Model (d) Full-context Model (Small-sized) (c) Previleged agent with $p(\text{STM}\rightarrow\text{LTM})=100\%$.
+In Figure 14, We display trajectories generated by different policies on three randomly selected $35\times 35$ tasks.
+Interestingly, we find that the main bottleneck of performance might arise from a lack of efficient exploration.
+Models with less context tend to explore less area.
+For instance, the random policy can only traverse a very small part of the maze.
+This observation could also explain the extraordinary challenge posed by the tasks: Traditional randomized exploration, a common technique in reinforcement learning, proves inadequate because it often revisits the same areas repeatedly without retaining information about which parts have been explored, thereby failing to cover broader regions effectively.
+Therefore, the agent must not only learn how to memorize paths and navigate but also how to explore more effectively, generating high-quality “training data” within its context to improve its future performance.
+Figure 15: Cases regarding the in-context improvement of world modeling (with $15\times 15$ mazes, small-sized full-context causal transformer).
+Each clip predicts $9$ steps into the future, conditioned on a fixed sequence of actions, and based on varying lengths of context denoted by $t$.
+Each clip consists of two sections: the upper section displays the ground truth observations from the current step to $9$ steps ahead (from left to right), while the bottom section shows the predicted future states up to $9$ steps ahead.
+In Figure 15, we illustrate a qualitative evaluation of the full-context causal transformer’s world modeling capabilities.
+Across the majority of instances investigated, a noticeable enhancement in future predicting is evident as the context length grows, suggesting that the model progressively adapts to the intricacies of the maze.
+This observation indicates the promising applicability of GPICL to the realm of adaptive embodied intelligence.
+A.5 Additional Diversity in the Maze World
+(a) 2D discrete
+(b) SURVIVAL tasks
+(c) narrow
+(d) wide
+(e) low
+(f) high
+(g) sparse
+(h) dense
+(i) close-sighted
+(j) far-sighted
+Figure 16: Additional variants by provided configurations in Maze World.
+It is worth noting that Maze World is a lightweight and fast engine with an easy-to-use Python API based on GYM [Brockman et al., 2016] for simulating navigation in various mazes.
+It runs smoothly at over 20 frames per second on a standard CPU.
+In addition to the aforementioned details, Maze World incorporates various configurations and hyper-parameters that can be adjusted to create a wide range of tasks.
+For example, it offers different types of observation spaces and action spaces, including discrete and continuous actions, as well as 2D and 3D observations (Figure 16a).
+Beyond navigation tasks, we introduce more challenging survival tasks, where the rewards provided by the PNTs are static within an episode but initially unknown (can be negative, Figure 16b).
+The basic environments can also be modified in terms of the size of the basic cell (Figure 16c,16d), the height of the ceiling (Figure 16e,16f), the density of obstacles (Figure 16g,16h), and the range of view (Figure 16i,16j).
+A.6 Details of Task Configurations
+In Table 2, we list the detailed settings for the tasks used in this paper.
+The rewards for reaching the target are by default related to the size of the maze.
+Additionally, we note that the density of obstacles controls the number of loops in the environments.
+We recommend using a density between 0.15 and 0.40.
+A density over 0.50 results in mazes without any loops, significantly reducing the difficulty of traversing the maze (or performing exploration), as an agent can easily traverse the entire world by always going right (or left).
+We believe this might be a crucial reason why, in Morad et al.
+[2023], where the agent is only required to navigate to a single target, context-free models can outperform memory-augmented models.
+Table 2: Setting of task configurations used in this paper
+Meta-Langauge
+Embedding size of the randomized langauge generator 32
+Hidden size of the randomized language generator 64
+Normalize factor for softmax of the randomized language generator ($\lambda$) 5
+Number of the hidden parameters ($|\theta|$, $n=2$) $7$K
+Number of the hidden parameters ($|\theta|$, $n=4$) $11$K
+Number of the hidden parameters ($|\theta|$, $n=8$) $19$K
+Maze Tasks
+Density of obstacles $0.36$
+Number of Potential Navigation Targets (PNTs) $10$
+Resolution of observation $128\times 128$
+Range of view 12.0 (default)
+Size of basic cell 2.0 (default)
+Rewards of arriving target ($15\times 15$) 0.57 (default)
+Rewards of arriving target ($25\times 25$) 1.24 (default)
+Rewards of arriving target ($35\times 35$) 2.06 (default)
+Height of ceiling 3.2 (default)
+Height of agents’ view points 1.6 (default)

### v6

@@
 In-context learning (ICL) empowers generative models to address new tasks effectively and efficiently on the fly, without relying on any artificially crafted optimization techniques.
-In this paper, we study extending ICL to address a broader range of tasks with an extended learning horizon and higher improvement potential, namely General-Purpose In-Context Learning (GPICL).
+In this paper, we study extending ICL to address a broader range of tasks with an extended learning horizon and higher improvement potential, namely General Purpose In-Context Learning (GPICL).
 To this end, we introduce two lightweight benchmarks specifically crafted to train and evaluate GPICL functionalities.
-Each benchmark encompasses a vast number of tasks characterized by significant task variance, facilitating meta-training that minimizes inductive bias.
-These tasks are also crafted to promote long-horizon in-context learning through continuous generation and interaction.
-These characteristics necessitate the models to leverage contexts and history interactions to enhance their capabilities, across domains such as language modeling, decision-making, and world modeling.
-Our experiments on the baseline models demonstrate that meta-training with minimal inductive bias and ICL from the ground up is feasible across all the domains we’ve discussed.
+Each benchmark encompasses a vast number of tasks characterized by significant task variance.
+These tasks are also crafted to promote long-horizon in-context learning through continuous generation and interaction, covering domains such as language modeling, decision-making, and world modeling.
+The benchmarks necessitate the models to leverage contexts and history interactions to enhance their capabilities, which we believe to be the key characteristics of GPICL.
+Our experiments indicate that the diversity of training tasks is positively correlated with the ability to generalize with ICL, but inversely correlated with zero-shot capabilities.
 Additionally, our findings indicate that the scale of parameters alone may not be crucial for ICL or GPICL, suggesting alternative approaches such as increasing the scale of contexts and memory states.
-00footnotetext: Works done at Baidu Inc.
-and University of Science and Technology of China.00footnotetext: Corresponds with: Fan Wang (fanwang.px@gmail.com) and Yang Cao (forrest@ustc.edu.cn)00footnotetext: Source code available at https://github.com/FutureAGI/L3C
+00footnotetext: source code will be available in the final version
 1 Introduction
@@
 The ICL capability allows the agent to handle various unseen tasks without modifying the model parameters, only resulting in changes to the hidden states and memories.
-Currently, the ICL capabilities of large language models are confined to tasks that solely involve interpreting natural language tokens and following instructions (i.e., mimicking demonstrations).
+Currently, the ICL capabilities of large language models are confined to natural language-based tasks and relatively naive tasks such as following instructions and mimicking demonstrations.
 However, they generally lack the ability to continually interact with real-world environments and engage in reinforcement learning from contexts.
@@
 GPICL Low High Long
-Figure 1: Conceptual diagram of in-context learning (ICL) and general-purpose in-context learning (GPICL)
-There are significant gap between artificial intelligence and biological intelligence, particularly regarding the functionality to continuously learn throughout their lifespans (Parisi et al., 2019).
-The enormous lifelong learning and adaptation potential of biological neural networks, forming abrupt contrast with their limited innate abilities at the initial stages (e.g., the stage of mammalian infancy), are also used to guide the design of machine intelligence systems (Zador, 2019; Wang et al., 2022; Schmidgall et al., 2024).
-A promising area in this line of research is general-purpose in-context learning (GPICL) (Kirsch et al., 2022; 2023).
-GPICL involves meta-training across a wide range of task classes and learning to interpret context within a variety of highly diverse tasks.
+Figure 1: Conceptual diagram of in-context learning (ICL) and general purpose in-context learning (GPICL)
+On the other hand, there are significant gap between artificial intelligence and biological intelligence, particularly regarding the functionality to continuously learn throughout their lifespans (Parisi et al., 2019).
+The enormous lifelong learning and adaptation potential of biological neural networks that forms abrupt contrast with their limited innate abilities at the initial stages (e.g., the stage of mammalian infancy) are also used to inspire the design of machine intelligence systems (Zador, 2019; Wang et al., 2022; Schmidgall et al., 2024).
+Among those a promising topic is general purpose in-context learning (GPICL) (Kirsch et al., 2022; 2023).
+GPICL involves meta-training across a wide range of task classes and learning to interpret in context to generalize to new tasks.
 However, the characteristic of diversity in this context can be challenging to quantify.
 We aim to enrich the concept of GPICL by incorporating additional features, as shown in Figure 1.
-Consider an agent that ideally utilizes all available data, including meta-training data and contextual information; The GPICL system should corresponds to low zero-shot capability, high ICL potential, and an extended ICL horizon.
+We consider the meta-learning framework in which a model is meta-trained in the outer loop to utilize contextual information for task adaptation in the inner loop.
+We identify a model with GPICL capability by three main characteristics: 1.
+Low zero-shot performance; 2.
+High ICL potential; 3.
+A long ICL horizon.
 (a) LLM mainstream learning pipeline
@@
 Figure 2: Learning pipelines based on GPICL versus the current LLM learning pipeline (Ouyang et al., 2022).
-Importance of GPICL.
-The objective of extending the in-context learning (ICL) horizon and potential is widely acknowledged.
-However, suppressing zero-shot capability seem to contradict the prevailing trend in large language models.
-To elucidate its importance further, we emphasize three key points:
-First, theoretically, the lower zero-shot capability in meta-training indicates a reduced reliance on inductive biases, thereby demonstrating sufficient diversity of meta-training tasks.
-As a result, the meta-training process focus primarily on the functionality of ICL instead of task-specific knowledge.
-Conversely, the attainment of zero-shot capability might suggest that the tasks possess a common knowledge base.
-While it could lead to faster adaption within the scope of these tasks, it might potentially restrict the model’s applicability beyond the scope of these tasks.
-Second, practically, meta-training the learning functionality requires an enormous number of tasks, which is challenging to collect in the real world.
-The use of abundant synthetic data in pre-training and meta-training can be essential, but tend to be non-factual and biased, presenting challenges such as the sim-to-real gap.
-Ideally, the meta-training stage of GPICL results in the loss of much task-specific knowledge while preserving the ICL functionality, thus helps prevent the model from the biases in the synthetic datasets.
-Building on this point, an alternative approach could be developed to replace the current mainstream learning pipeline of Pre-training & Fine-tuning & ICL with Meta-training & GPICL (Figure 2).
-In the new pipeline, initially, massive synthesized data, potentially containing a wealth of nonfactual and biased information, is used exclusively to cultivate the GPICL functionality.
-Subsequently, a considerable amount of high-quality, unbiased, well-aligned, real-world data is employed to enrich the contexts and memory, enabling the agent to achieve capabilities approaching those of humans.
-Third, since a model’s performance typically depends on both in-context learning (ICL) and zero-shot generalization capabilities, it is challenging to investigate and improve ICL independently.
-Isolating the ICL capability alone would be beneficial for analysis on model architectures and learning algorithms.
-Guidelines for designing GPICL benchmarks.
-We note that a major impediment to research in this area is the lack of benchmarks that adhere to the GPICL standard.
-We suggest that those benchmarks should satisfy the following standard:
-  • Vast number of variant tasks with minimal inductive bias: It is desirable to ensure that the meta-training and meta-testing tasks are sufficiently diverse, preventing the transfer of task-specific knowledge among different tasks.
-This guarantees that the model’s performance is primarily driven by its in-context learning capabilities rather than inductive biases.
-It is also essential to require numerous tasks to prevent the model from hacking in-context learning (ICL) by merely memorizing all possible tasks and performing task identification, a limitation that many meta-learning benchmarks fall into when given limited tasks, as is discussed in Kirsch et al.
+We propose two lightweight benchmarks that adhere to the GPICL standard and are capable of generating unlimited synthetic data: 1.
+Meta-Language: This benchmark generates random yet consistent language patterns for meta-training the ICL functionality, enabling it to master a novel language from scratch.
+2.
+Maze World: This benchmark consists of a variety of mazes with different targets that are randomly generated for navigation purposes.
+It is designed for training and evaluating models capable of navigating in various scenarios through observation, mapping, localization, and planning.
+We provide experiments of baseline models composed of language modeling, world modeling, and decision modeling based on the Transformer backbone and its variants, as described in Vaswani et al.
 (2017).
-  • Interactive and generative tasks: GPICL should include task exploration and completion processes that are both interactive and generative, to ensure obtaining complex functionalities such as in-context reinforcement learning.
-Consequently, benchmarks that solely focus on encoding a single sample characterized by extended contexts and features and conducting single-step predictions or classifications do not meet the criteria.
-  • Lifelong in-context learning: To guarantee the capability of adapting complex tasks, GPICL should handle very long-term dependencies that can scale up to millions or even billions of steps, effectively supporting the learning process over extensive periods (ICL horizon $T$ should be “from cradle to grave”).
-It is important to clarify that $T$ should denote the “minimum steps” possible given “ideal ICL functionality”, excluding scenarios requiring extensive contexts that are primarily due to inefficient learning functionalities.
-Contribution of this paper: We propose two critical benchmarks that adhere to the GPICL standard and are capable of generating unlimited synthetic data: 1.
-Meta-Language: We create a randomized yet consistent language pattern to meta-train models, demonstrating that a model, without prior exposure to natural language corpora, can acquire the ability to learn a new language from scratch through ICL; 2.
-Maze World: We generate an unlimited number of tasks within randomized mazes, where the model is meta-trained solely through context to learn exploration and navigation.
-For both benchmarks, we provide limited experiment results of baseline models based on Transformer (Vaswani et al., 2017) backbone and its variants.
-The primary goal of this paper is neither to introduce a new model or learning methodology, nor to provide an exhaustive survey of the existing ICL systems; rather, we aim to demonstrate through these baselines the potential for these benchmarks to benefit the research of ICL and GPICL.
-One crucial conclusion from these limited experiments may be that the scale of parameters alone is not crucial for ICL or GPICL.
-This finding aligns with previous research indicating that the size of contexts, and the scale of the memory states can play a more significant role (Kirsch et al., 2022; Wang et al., 2022) in learning from contexts.
+It is shown that: 1.
+With auto-regressive models, both benchmarks exhibit low zero-shot performance and high potential for generalization with ICL; 2.
+Increasing the diversity of meta-training tasks can naturally lead to decreased zero-shot performance, improved generalization to unseen tasks, and an extended ICL horizon, implying that sufficiently diverse tasks can lead to effective GPICL; 3.
+The performance of GPICL is not strongly correlated with the scale of parameters once a certain threshold is exceeded.
+On the other hand, the size of the context significantly influences the model’s performance, which aligns with the findings in Kirsch et al.
+(2022); Wang et al.
+(2022).
+The results strongly suggest that, in evaluations, the position-wise performance of the auto-regressive model should be investigated in detail, rather than averaging the performance across different context lengths.
+Our findings reveal a potential alternative approach to replace the current mainstream learning pipeline of large-scale language models, as shown in Figure 2.
+Typical Large Language Models (LLMs) go through three stages: Pre-training, Fine-tuning, and ICL to train and customize.
+With GPICL, it is possible to simplify this pipeline to just two stages: Meta-training and GPICL.
+Moreover, given the data efficiency of GPICL compared with gradient-based methods, it is possible to base meta-training on low-quality synthetic data and use high-quality data more efficiently to master novel skills.
 2 Learning to learn an randomized language: Meta-Language
@@
 2.3.1 Towards Meta-Language Model
-Figure 3: Perplexity per position on evaluation data sampled by different language complexity.
+Figure 3: Comparing evaluations on meta-language sampled by different language complexity.
+Horizontal axis represent the position or context length, and vertical axis represent the perplexity (the lower the better).
 We demonstrate that this pseudo-language trained model can be regarded as a meta-language model designed to adapt to any conceivable language through GPICL.
@@
 We believe research in this area (Tay et al., 2022) could benefit from this benchmark.
-Figure 4: Evaluating language models of various sizes on Meta-Language benchmark with differing levels of complexity ($n$).
+Figure 4: Comparing performances of language models of different sizes on the the same test set.
+Horizontal axis represent the position or context length, and vertical axis represent the perplexity (the lower the better).
 3 Learning to navigate an randomized world: Maze World
@@
 3.3 Baseline Performances
-(a) $\mathcal{L}_{wm}(t)$
-(b) $\mathcal{L}_{pm}(t)$
 Figure 7: Static assessment of the ability to imitate previleged agents ($\mathcal{L}_{pm}$) and predict the next frame ($\mathcal{L}_{wm}$) across each time step ($t$).
@@
 These results may indicate that imitating the exploration strategy is more challenging.
-(a) Maze $15\times 15$
-(b) Maze $25\times 25$
-(c) Maze $35\times 35$
 Figure 8: Interactive evaluation results of the randomized, previleged and casual-modeling based agents, depicted by plotting accumulated rewards against simulated time steps.
@@
 We believe it is beneficial to incorporate on-policy trajectories in the dataset in the future, either in reinforcement learning or DAgger (Ross et al., 2011).
-(a) Maze $15\times 15$
-(b) Maze $25\times 25$
-(c) Maze $35\times 35$
 Figure 9: Results regarding the in-context improvement of world modeling in Maze World.
-We examine the compound error associated with a 9-step auto-regressive future predicting process, using the learned world models that differ in the context length ($t\in\{1,100,1000,2000\}$).
-The figures presented compare different models to the context-free world model as the baseline.
-The horizontal axis measures the number of imagined steps, with each curve illustrating a distinct context length.
-The mean square error is averaged across 64 tasks, and the shaded area provides a 95% confidence interval for these measurements.
+We assess the mean square error (MSE) for forecasting subsequent frames.
+When $k=1$, it indicates the error associated with predicting the immediately following frame.
+When $k=4$, it denotes the cumulative error for predicting 4 frames ahead in an autoregressive manner.
+The predictions are initiated from various context lengths (or time step), with $t$ taking values from the set {1, 100, 1000, 2000}.
+The mean square error is averaged across 64 procedural generated tasks, and the shaded area provides a 95% confidence interval for these measurements.
 In-Context Improvement of World Modeling.
-Figure 9 explores the performance of world models in interactive evaluations.
-At steps $1$, $100$, $1000$, and $2000$ of each task, the world model is tasked with “imagining” $9$ steps into the future.
-For the smaller $15\times 15$ mazes, we observe a significant improvement in world modeling as context length increases from $1$ to $1000$, evidenced by a continuous decrease in prediction error.
-However, this improvement plateaus between $1000$ and $2000$ steps.
-With sufficient context, the cumulative error over 9 steps remains lower than that of an average guess, which is the performance of the context-free model.
-For larger mazes of $25\times 25$ and beyond, and when predicting more than 5 steps into the future, none of the trained models outperform an average guess.
-Additionally, the partial-context model does not enhance its predictive capabilities as context length increases, which aligns with expectations.
-The results indicate that the proposed model demonstrates a reasonable capacity to utilize contextual information for future predictions, yet there is considerable room for improvement, especially in solving larger-scale mazes and the efficient utilization of long-term dependencies.
+In Figure 9, the world model is evaluated by comparing the ground truth future frames and predicted future frames at steps $1$, $100$, $1000$, and $2000$ of each interactive run.
+We demonstrate the performance of world models in predicting both the immediate future ($k=1$) and a slightly more distant future ($k=4$).
+The latter is more reliant on long-term memory, making it inherently more difficult.
+In the case of smaller $15\times 15$ mazes, we notice a substantial enhancement in world modeling as the context length expands from $1$ to $1000$, which is validated by a consistent reduction of prediction errors for both $k=1$ and $k=4$.
+For larger-scale mazes, the challenge of predicting $k=4$ is amplified, as retaining the entire map in memory becomes increasingly challenging.
+The partial-context model does not enhance its predictive capabilities as context length increases over $100$, which aligns with expectations that it can not keep a long-term memory.
+The results indicate that the full-context causal transformer demonstrates a reasonable capacity to utilize contextual information for world modeling, yet there is considerable room for improvement, especially in solving larger-scale mazes and the efficient utilization of long-term dependencies.
 For further insights and cases, refer to Appendix A.4.
+4 Emergence of Generalizability
+Figure 10: Performance of the meta-language model trained with 500,000 sequences generated from varying numbers of pre-selected tasks.
+The vertical axis represents the average perplexity in evaluation (lower values indicate better performance), and the horizontal axis represents the context length.
+Each point in the seen tasks evaluation is the average of 4,000 sequences generated from the pre-selected tasks.
+Each point in the unseen tasks evaluation is based on the PG-19 test set.
+Figure 11: Performance of the causal transformer model trained with 64,000 Maze World sequences generated from varying numbers of pre-defined tasks.
+The left vertical axis represents the accumulated reward (higher values indicate better performance), while the right vertical axis represents the prediction error of world modeling in the interactive evaluation (lower values indicate better performance).
+The horizontal axis represents the context length or time steps.
+For the prediction error of world modeling, we evaluated the mean squared error (MSE) for predicting 1 and 4 steps into the future at context lengths of 1, 100, 1000, and 2000.
+Each evaluation point is the average of 64 runs, with the shaded area representing the 95% confidence interval.
+To illustrate the relationship between the two benchmarks and GPICL, we delve into how the ICL capability is linked to the diversity of meta-training data in these benchmarks.
+In this section, we pre-select a set of tasks to generate the sequences (dataset) for meta-training.
+Unlike the procedural generation of tasks for each sequence, this approach reduces the diversity of both the task-set and the dataset, by allowing multiple sequences to be derived from a single task.
+For example, in Meta-Language, each task is defined by the parameters of the n-gram generator ($\theta$), allowing for the generation of different sequences from a single task through softmax sampling.
+Similarly, in Maze World, each task is characterized by the layout, textures, and the position of the PNTs (potential navigation targets), with different trajectories being generated by sampling behavior policies.
+To examine the effect of task diversity on ICL capability, we create datasets of the same size but with varying numbers of pre-defined tasks.
+In Meta-Language, we generate 500,000 sequences (2 billion tokens) using 10, 1,000, and 100,000 pre-sampled tasks, respectively.
+In Maze-World, we synthesize 64,000 sequences (126 million frames) with 1, 10, 100, and 10,000 pre-sampled tasks for the $15\times 15$ mazes.
+Given that the performance of standard-sized models and small-sized models was nearly indistinguishable in our previous experiments, we use only small-sized models for both benchmarks.
+Each meta-trained model is assessed on both ’seen’ tasks (those used to generate the meta-training data) and ’unseen’ tasks.
+The outcomes are displayed in Figures 10 and 11.
+Several conclusions have been validated across both benchmarks: First, we observe a continuously increasing asymptotic performance (the performance with sufficiently long context) in unseen tasks as the number of meta-training tasks increases, a phenomenon also noted by Kirsch and Schmidhuber (2021).
+This finding confirms that the capability for task generalization is enhanced.
+Second, which is less discussed in the previous work, for seen tasks, performance does not consistently improve with the increase in the number of tasks; in contrast, it declines in most of the groups.
+Additionally, we notice an expansion of the ICL horizon, identified by the length of context during the performance improvement phase, and a reduction in zero-shot performance, identified by the level of performance at the first time step (where context length is 0), in both seen and unseen tasks.
+We hypothesize that this may be a consequence of the model transitioning from task-identification to learning-to-learn.
+In this process, the model retains less specific information about each task but gains improved generalizability and the ability to learn through ICL.
+The decline in performance for seen tasks could also be attributed to the insufficient context length we employ (up to 2,048 and 4,096) and the size of the dataset.
+Last, we observe a slower convergence in meta-training as the number of tasks increases, despite the dataset size remaining constant (Figure 15).
+This suggests that the process of learning to learn is more challenging than mastering specific tasks.
+It is important to note that the observed increase in generalizability, the expansion of the ICL horizon, and the reduction in zero-shot capability are all in line with the definition of GPICL that we previously mentioned.
 5 Related Work
@@
 This work can be expanded along two directions: 1.
-The development of more realistic GPICL benchmarks, for instance, by adding greater complexity to the meta-language; 2.
+The development of more realistic GPICL benchmarks, for instance, by including additional complexity to the meta-language; 2.
 Utilizing the benchmarks to uncover more sophisticated models (including memory-augmented models) and optimization techniques (such as reinforcement learning) to further extend the GPICL horizon and enhance performance gains.
@@
 Notably, within the Maze World framework, each temporal increment is encoded by a pair of tokens: an observation and its corresponding action.
-For the optimization of the Meta-Language and Maze World baseline models, we employed the Noam decay scheduler [Vaswani et al., 2017].
+For the optimization of the Meta-Language and Maze World baseline models, we employed the Noam decay scheduler [49].
 This scheduler peaks at a learning rate of $10^{-3}$ following an initial warm-up phase spanning $1,000$ steps.
-For the image encoder (11M parameters) and image decoder (14M parameters), we employ convolutional and deconvolutional layers augmented with ResNet blocks [He et al., 2016].
+For the image encoder (11M parameters) and image decoder (14M parameters), we employ convolutional and deconvolutional layers augmented with ResNet blocks [14].
 The detailed structure of the VAE is depicted in Figure 12.
@@
 Figure 13: Variant attention masks of causal modeling
-(a) Meta-Language Model
-(b) World Modeling
-(c) Policy Modeling
-(d) World Modeling
-(e) Policy Modeling
-Figure 12: The training loss with respect to the steps of optimization. Notice that the partial-context and context-free causal transformers are initialized from the trained full-context transformers.
-Figure 12 presents the training loss versus the number of iterations for the Meta-Language and Maze World baseline models.
+(a) $-logP$ (Meta-Language-Model)
+(b) $\mathcal{L}_{wm}$ (full-context)
+(c) $\mathcal{L}_{pm}$ (full-context)
+(d) $\mathcal{L}_{wm}$ (partial-context))
+(e) $\mathcal{L}_{pm}$ (partial-context)
+Figure 14: The training loss with respect to the steps of optimization in training with procedural generated tasks and sequences.
+Notice that the partial-context and context-free causal transformers are initialized from the trained full-context transformers.
+(a) $-logP$ (Small-sized)
+(b) $\mathcal{L}_{wm}$ (Small-sized)
+(c) $\mathcal{L}_{pm}$ (Small-sized)
+Figure 15: The training loss with respect to the steps of optimization in training with sequences generated from pre-selected tasks.
+Figure 14 presents the training loss versus the number of iterations for training the baseline in section 2 and A.5.
 During training baselines for Meta-Language tasks, we observed that standard-sized models converge relatively slowly on the training set.
@@
 The context-free and partial-context causal transformers are initialized from the full-context causal transformer, and the hyperparameters remain unchanged.
+Figure 15 presents the training loss versus the number of iterations for the experiments in section 4. Both group are experimented on small-sized models.
 A.2 Data Collection Strategy for Maze World
-For dataset collection and static evaluation of Maze World, inspired by dataset aggregation [Ross et al., 2011] and noise distillation [Zisman et al., 2023], we utilize distinct behavior and reference policies.
+For dataset collection and static evaluation of Maze World, inspired by dataset aggregation [37] and noise distillation [55], we utilize distinct behavior and reference policies.
 The behavior policy is determined by previleged agent where $p(\text{STM}\rightarrow\text{LTM})$ randomly sampled between $[0,50\%]$.
@@
 The results, depicted in Fig.
-13, show that MetaLM is capable of learning to memorize and correct both English vocabulary and mathematical equations, validating that its ICL capability is independent of vocabulary settings.
+16, show that MetaLM is capable of learning to memorize and correct both English vocabulary and mathematical equations, validating that its ICL capability is independent of vocabulary settings.
 Furthermore, the longer the context provided, the fewer errors appeared in the outputs, indicating a clear trend of improvement.
@@
 Figure 19: Additional variants by provided configurations in Maze World.
-It is worth noting that Maze World is a lightweight and fast engine with an easy-to-use Python API based on GYM [Brockman et al., 2016] for simulating navigation in various mazes.
+It is worth noting that Maze World is a lightweight and fast engine with an easy-to-use Python API based on GYM [2] for simulating navigation in various mazes.
 It runs smoothly at over 20 frames per second on a standard CPU.
@@
 Beyond navigation tasks, we introduce more challenging survival tasks, where the rewards provided by the PNTs are static within an episode but initially unknown (can be negative, Figure 19b).
-The basic environments can also be modified in terms of the size of the basic cell (Figure 16c,16d), the height of the ceiling (Figure 16e,16f), the density of obstacles (Figure 16g,16h), and the range of view (Figure 16i,16j).
+The basic environments can also be modified in terms of the size of the basic cell (Figure 19c,19d), the height of the ceiling (Figure 19e,19f), the density of obstacles (Figure 19g,19h), and the range of view (Figure 19i,19j).
 A.6 Details of Task Configurations
@@
 A density over 0.50 results in mazes without any loops, significantly reducing the difficulty of traversing the maze (or performing exploration), as an agent can easily traverse the entire world by always going right (or left).
-We believe this might be a crucial reason why, in Morad et al.
-[2023], where the agent is only required to navigate to a single target, context-free models can outperform memory-augmented models.
+We believe this might be a crucial reason why, in [29], where the agent is only required to navigate to a single target, context-free models can outperform memory-augmented models.
 Table 2: Setting of task configurations used in this paper